### v2

# Ramified Approximation and Semistable Reduction

Abstract
Let $K$ be a complete discretely valued field. An extension $L/K$ is weakly totally ramified if the residue extension is purely inseparable. We sharpen a result of Ax by showing that any Galois-invariant disk in the algebraic closure of $K$ contains an element that generates a separable weakly totally ramified extension. As an application, we prove that elliptic curves and dynamical systems on $\mathbb{P}^{1}$ achieve semistable reduction over a separable weakly totally ramified extension of the base field. We also obtain several arithmetic consequences for torsion points on elliptic curves and preperiodic points for dynamical systems.
1 Introduction
We set the following notation for this article:
$K$ complete discretely valued field
$v$ the discrete valuation on $K$, normalized so that $v(K^{\times})=\mathbb{Z}$
$\tilde{K}$ residue field of $K$
$K_{\mathrm{a}}$ fixed algebraic closure of $K$
$\mathbb{C}_{K}$ completion of $K_{\mathrm{a}}$ with respect to the unique extension of $v$
$G_{K}$ $\operatorname{Aut}_{K}(K_{\mathrm{a}})=\operatorname{Aut}_{K}(\mathbb{C}_{K})$, the group of field automorphisms fixing $K$
For $a\in\mathbb{C}_{K}$ and $r\in\mathbb{R}\cup\{\infty\}$, define
$D(a,r)=\{b\in\mathbb{C}_{K}\colon v(b-a)\geq r\}$
to be the disk with center $a$ and radius $r$. The group $G_{K}$ maps disks to disks, and a disk $D$ is $G_{K}$-invariant if and only if there is an irreducible polynomial $f\in K[z]$ whose roots belong to $D$. Given a $G_{K}$-invariant disk, what can be said about the minimum-degree algebraic elements in it? A first answer was given by Ax in 1970:
Ax’s Lemma ([1, Lemma 2]).
Let $f\in K[z]$ be a nonconstant polynomial of degree $d=qm$, where $q$ and $m$ are coprime, and where $q=1$ if $\mathrm{char}(\tilde{K})=0$ and $q$ is a power of $\mathrm{char}(\tilde{K})$ otherwise. Assume that $D$ is a disk containing the roots of $f$. There exists $\alpha\in D$ such that $[K(\alpha):K]\leq q$.
Following work of Tate [23, §3.3], Ax was interested in determining the $G_{K}$-invariant transcendental elements of $\mathbb{C}_{K}$. And while Ax’s Lemma implies that a minimum-degree element in a $G_{K}$-invariant disk generates a $p$-power extension of $K$, its proof tells us nothing about the residue extension of the field generated by an element of minimum degree. Our main result will remedy this defect.
We say that a finite extension $L/K$ is weakly totally ramified if its residue extension is purely inseparable. The properties “totally ramified” and “weakly totally ramified” agree when the residue field of $K$ is perfect, so little is lost on a first reading by conflating these notions. We say that $L/K$ is weakly totally wildly ramified if it is weakly totally ramified and either $L=K$ or the ramification index is a power of the residue characteristic of $K$. (See Section 2 for more on these definitions.)
Ramified Approximation.
Let $\beta\in K_{\mathrm{a}}$, and suppose that $D$ is a $G_{K}$-invariant disk containing $\beta$. There exists $\alpha\in D\cap K_{\mathrm{a}}$ such that $K(\alpha)/K$ is separable and weakly totally wildly ramified. Moreover, the following hold:
  • (Inclusion of residue fields) We may take $\tilde{K}\subset\widetilde{K(\alpha)}\subset\widetilde{K(\beta)}$.
  • (Inclusion of value groups) We may take $v(K^{\times})\subset v\big{(}K(\alpha)^{\times}\big{)}\subset v\big{(}K(\beta)^{\times}\big{)}$.
In particular, if $K$ has residue characteristic zero, then $\alpha\in K$. If $K$ has residue characteristic $p>0$, then $[K(\alpha):K]$ is a power of $p$ that divides $[K(\beta):K]$.
The proof uses MacLane’s method of approximants [8, 9] to construct a sequence of semivaluations $V_{1},\ldots,V_{n}$ on $K[z]$ satisfying $V_{1}(f)<V_{2}(f)<\cdots<V_{n}(f)=\infty$, where $f$ is the minimal polynomial of $\beta$. The first $V_{j}$ whose center lies in $D$ gives rise to an element $\alpha$ with all of the desired properties. As a result, the extension $K(\alpha)/K$ may be computed explicitly provided we have a model for computation in the valued field $K$ and a means to factor polynomials over the residue field. For example, this is feasible if $K$ is a finite extension of the $p$-adic field $\mathbb{Q}_{p}$ or of the function field $\mathbb{F}_{p}(T)$. Other recent applications of MacLane’s method to geometric questions are given in [12, 13]; both of these articles address ramification questions, although of a different flavor than those in the present study.
Remark 1.1.
Though the statement of Ramified Approximation does give inclusions of local invariants, it is typically not true that we have $K(\alpha)\subset K(\beta)$. For example, take $K=\mathbb{Q}_{2}$ and $v=\operatorname{ord}_{2}$. If $\beta$ is given by
$\beta^{4}+2\beta^{3}+4\beta^{2}+12\beta+12=0,$
then the proof produces an element $\alpha$ satisfying $\alpha^{2}+2=0$. One verifies that $D=D(\beta,3/4)$ is the minimum disk containing all conjugates of $\beta$, that $\alpha\in D$, but that $K(\alpha)\not\subset K(\beta)$.
Suppose the residue characteristic of $K$ is $p>0$. The residue degree and ramification index of a weakly totally wildly ramified extension are both powers of $p$. This means we can apply Ramified Approximation to the irreducible factors of an arbitrary nonconstant polynomial in order to sharpen Ax’s Lemma, thus proving a Conjecture of Benedetto [3, Conj. 2]:
Corollary 1.2.
Let $f\in K[z]$ be a nonconstant polynomial of degree $d=qm$, where $q$ and $m$ are coprime, and where $q=1$ if $\mathrm{char}(\tilde{K})=0$ and $q$ is a power of $\mathrm{char}(\tilde{K})$ otherwise. Assume that $D$ is a disk of finite radius containing the roots of $f$. There exists $\alpha\in D$ such that $[K(\alpha):K]\leq q$ and $K(\alpha)/K$ is separable and weakly totally wildly ramified.
We now turn to a characterization of the minimum-degree elements in $G_{K}$-invariant disks. The question is settled if $K$ has residue characteristic zero: any such disk contains an element of $K$. The next result is an immediate consequence of Ramified Approximation.
Corollary 1.3.
Let $K$ be a complete discretely valued field of residue characteristic $p>0$. Suppose that $D$ is a $G_{K}$-invariant disk.
  1. (a)
    Let $L/K$ be a field extension of minimum degree such that $D\cap L\neq\varnothing$. Then $L/K$ is a weakly totally wildly ramified extension.
  2. (b)
    Let $M/K$ be a field extension of minimum ramification index such that $D\cap M\neq\varnothing$. The ramification index of $M/K$ is $p^{m}$ for some $m\geq 0$, and there exists a weakly totally wildly ramified extension $M^{\prime}/K$ with ramification index $p^{m}$ such that $D\cap M^{\prime}\neq\varnothing$.
  3. (c)
    Let $N/K$ be a field extension of minimum residue degree such that $D\cap N\neq\varnothing$. The residue degree of $N/K$ is $p^{n}$ for some $n\geq 0$, and there exists a weakly totally wildly ramified extension $N^{\prime}/K$ with residue degree $p^{n}$ such that $D\cap N^{\prime}\neq\varnothing$.
Corollary 1.3(b) sharpens a result of Rozensztajn when $p>2$ [14, Theorem 1.1.1]. Consequently, it has implications for computing the complexity of standard subsets of $\mathbb{P}^{1}(K_{\mathrm{a}})$ that arise in the description of 2-dimensional crystalline representations of $G_{\mathbb{Q}_{p}}$.
We now turn to an application of Ramified Approximation: weakly totally ramified extensions of the base field are sufficient to achieve semistable reduction for elliptic curves and dynamical systems on $\mathbb{P}^{1}$. The criterion of Néron, Ogg, and Shafarevich suggests that ramification should be the only obstruction to semistable reduction: if $K$ has perfect residue field, then an elliptic curve $E_{/K}$ has good reduction if and only if its $\ell$-adic Tate module is unramified for $\ell\neq\mathrm{char}(\tilde{K})$ [20, VII.7]. Generalizations of this criterion for semistable reduction of abelian varieties were proved by Serre–Tate and Grothendieck [5, §7.4]. We are unaware of any results along these lines in the setting of arboreal representations of dynamical systems.
Theorem A.
Let $K$ be a complete discretely valued field, and let $E$ be an elliptic curve over $K$. There is a separable weakly totally ramified extension $L/K$ and a model for $E_{L}$ that admits semistable reduction. If $E$ has potential multiplicative reduction, then we may take $L/K$ to be at most quadratic. If $E$ has potential good reduction, then we may take $[L:K]$ to be
  • a proper divisor of $12$ if $\mathrm{char}(\tilde{K})\neq 2,3$;
  • a divisor of $12$ if $\mathrm{char}(\tilde{K})=3$; or
  • a divisor of $24$ if $\mathrm{char}(\tilde{K})=2$.
The novel feature of Theorem A is that we may take $L/K$ to be weakly totally ramified. This is quite easy to show when $\mathrm{char}(\tilde{K})\neq 2$ or $3$, a difficult exercise when $\mathrm{char}(\tilde{K})=3$, and requires the power of Ramified Approximation when $\mathrm{char}(\tilde{K})=2$. The proof gives an algorithm for finding the field extension $L/K$, and we suspect that it produces the minimal such extension.
The word “weakly” is necessary in Theorem A. For example, take $K=\mathbb{F}_{3}(T){\left(\!\left(\pi\right)\!\right)}$ to be the field of formal Laurent series over $\mathbb{F}_{3}(T)$ with the valuation $\operatorname{ord}_{\pi}$, and define
$E_{/K}\colon y^{2}=x^{3}+\pi x^{2}-\pi x-T.$
Then $E$ has integral $j$-invariant, and an inseparable extension of the residue field is required in order to $\pi$-adically separate the $2$-torsion and achieve good reduction. For more examples, including some in residue characteristic 2, see [22].
It is well known that if $m$ is coprime to the residue characteristic of $K$ and $E_{/K}$ is an elliptic curve with good reduction, then the reduction map is injective on $m$-torsion: $E(K)[m]\hookrightarrow E(\tilde{K})$. Since the residue field does not grow when we pass to a totally ramified extension, this observation allows us to make certain weak uniformity statements about torsion. The following is an example of such a statement:
Corollary 1.4.
Let $F$ be a number field, and let $E_{/F}$ be an elliptic curve. Suppose that $E$ has potential good reduction at places $\mathfrak{p}$ and $\mathfrak{q}$ of $F$ such that $\mathfrak{p}$ lies above 2 and is totally ramified, and $\mathfrak{q}$ lies above 3 and is totally ramified. Then
#$\#E(F)_{\mathrm{tors}}\in\{1,2,3,4,5,6,12\}.$
Proof.
By Theorem A, there is a totally ramified extension of the completion $F_{\mathfrak{p}}$ over which $E$ has good reduction. The residue field of such an extension is $\mathbb{F}_{2}$. An elliptic curve over $\mathbb{F}_{2}$ has at most 5 rational points, so the prime-to-$2$ part of $E(F)_{\mathrm{tors}}$ is of size $1,3$, or $5$. Similarly, there is a totally ramified extension of the completion $F_{\mathfrak{q}}$ over which $E$ has good reduction. An elliptic curve over $\mathbb{F}_{3}$ has at most 7 rational points, so the prime-to-$3$ part is of size $1,2,4,5$ or $7$. The only possible torsion orders that fit these numerics are the ones given in the statement of the corollary. ∎
A dynamical system $f\colon\mathbb{P}^{1}_{K}\to\mathbb{P}^{1}_{K}$ — i.e., a morphism of degree at least 2 — is said to have semistable reduction if its normalized homogeneous resultant has minimal valuation among all possible coordinate changes over $\mathbb{C}_{K}$. If this minimal valuation is 0, then $f$ has good reduction. There is an equivalent definition in terms of GIT semistability. (See §7 for references and more details.) Rumely proved that the dynamical system $f$ admits semistable reduction after a finite extension $L/K$ of degree at most $(\deg(f)+1)^{2}$, and that $L/K$ is either trivial or ramified [15, §3]. We strengthen both of these statements.
Write $p\geq 0$ for the residue characteristic of $K$. For $d\geq 2$, define
$q(d)=\begin{cases}p^{\operatorname{ord}_{p}(d)}&\text{ if $p>0$}\\ 1&\text{ if $p=0$}.\end{cases}$
Set $A(p,d)=(d+1)\max\left\{q(d+1),q(d-1)\right\}$, and set
$B(p,d)=\begin{cases}(d-1)\ q(d)&\text{ if $p>0$ and $d\equiv 0\pmod{p}$}\\ d\ q(d-1)&\text{ if $p>0$ and $d\equiv 1\pmod{p}$}\\ d+1&\text{ otherwise}.\end{cases}$
Note that $A(p,d)\leq(d+1)^{2}$, while $B(p,d)\leq d(d-1)$ if $d>2$. Both quantities are at most $d+1$ when $K$ has residue characteristic zero or when $d\not\equiv 0,\pm 1\pmod{p}$.
Theorem B.
Let $K$ be a complete discretely valued field with $\mathrm{char}(\tilde{K})=p\geq 0$, and let $f\colon\mathbb{P}^{1}_{K}\to\mathbb{P}^{1}_{K}$ be a morphism of degree $d\geq 2$. There is a separable weakly totally ramified extension $L/K$ over which $f$ attains semistable reduction. Moreover, we may take $[L:K]\leq A(p,d)$ in general, and we may take $[L:K]\leq B(p,d)$ if $f$ has potential good reduction.
Benedetto proved the numerical part of this result when $f$ has potential good reduction [3, Thm. A], and he conjectured that one can choose the extension $L/K$ to be totally ramified when $K$ has perfect residue field. Theorem B extends and proves this conjecture.
Similar to the case of elliptic curves, there are a number of weak uniformity results that can be obtained for periodic points of dynamical systems over number fields. Good reduction versions of these were stated by Morton and Silverman [10], and we can immediately extend all of them to the case of potential good reduction. For example:
Corollary 1.5.
Let $F$ be a number field and let $f\colon\mathbb{P}^{1}_{F}\to\mathbb{P}^{1}_{F}$ be a dynamical system of degree at least 2. Suppose that $P\in\mathbb{P}^{1}(F)$ is periodic with minimal period $n$.
  • (a)
    Let $\mathfrak{p}$ and $\mathfrak{q}$ be finite places of $F$ with distinct residue characteristics for which $f$ has potential good reduction. Then $n\leq(N_{F/\mathbb{Q}}\mathfrak{p}^{2}-1)(N_{F/\mathbb{Q}}\mathfrak{q}^{2}-1)$.
  • (b)
    If $F=\mathbb{Q}$ and $f$ has potential good reduction at $2$ and $3$, or at $3$ and $5$, then $n\mid 24$.
Proof.
For a given place of potential good reduction, we may apply Theorem B to pass to a totally ramified extension at which $f$ attains good reduction. Since the residue field is unaffected by such an extension, the proofs of [10, Cor. B and Cor. C] apply. ∎
The typical approach to obtaining a semistable model for an elliptic curve is to make a field extension that rationalizes certain torsion points [20, VII.5.4, A.1.4]. The proof of Theorem A uses a kind of “approximate rationalization” technique: it suffices to nearly rationalize torsion over a weakly totally ramified extension. The same idea works for dynamical systems: one needs to approximately rationalize certain periodic points. Ramified Approximation is the key ingredient behind this technique. However, it is a fundamentally 1-dimensional tool, and so it is unclear if it can be used to extend Theorems A and B to curves of genus at least 2, general abelian varieties, or dynamical systems on $\mathbb{P}^{n}_{K}$.
Some progress can be made if one places further restrictions on the field $K$. Suppose the residue field of $K$ is quasi-finite — i.e., $\tilde{K}$ admits a unique finite extension of every degree. Clark observed that an abelian variety over $K$ attains semistable reduction over a totally ramified extension of $K$ [7], though his argument does not control the degree of the extension. The argument would apply to dynamical systems $f\colon\mathbb{P}^{n}_{K}\to\mathbb{P}^{n}_{K}$ if we had an analogue of the Néron model that behaved well under unramified extensions.
Conjecture 1.6.
Let $K$ be a complete discretely valued field.
  • Let $A_{/K}$ be an abelian variety. There is a separable weakly totally ramified extension $L/K$ and a model for $A_{L}$ that admits semistable reduction.
  • Let $f\colon\mathbb{P}^{n}_{K}\to\mathbb{P}^{n}_{K}$ be a dynamical system of degree $d\geq 2$. There is a separable weakly totally ramified extension $L/K$ and a model for $f_{L}$ that admits semistable reduction.
As a final remark, we note that for many applications, the field $K$ will not be complete. Krasner’s Lemma can be used to deduce algebraic analogues of Ramified Approximation and Theorems A and B, with the same bounds for the degrees of the necessary extensions.
We recall or prove various properties of discretely valued fields in §2; additional notational conventions will also be laid out there. We then discuss semivaluations on $K[z]$ from two different vantages: as the points of the analytic affine line over $K$ in §3, and as inductive valuations in MacLane’s theory of approximants in §4. We turn to the proof of Ramified Approximation in §5. We prove Theorems A and B in §6 and §7, respectively.
Acknowledgments. While I was sure many years ago that Theorems A and B are true, they would not have been proved without the wealth of insight and encouragement I gained from talking with Andrew Obus, Rob Benedetto, Bob Rumely, and David Zelinsky. My sincerest appreciation goes to all of you.
2 Preliminaries on complete discretely valued fields
We begin by defining several concepts related to complete discretely valued fields and their extensions. These are well known when the residue field is perfect, but there is some fussiness over inseparable residue extensions in the imperfect case. Useful references include [11, Ch. II.6] and [18, Ch. I.4].
Let $K$ be a field that is complete with respect to a discrete valuation $v$. Without loss of generality, we may assume that $v(K^{\times})=\mathbb{Z}$. We write $\mathcal{O}_{K}$ for the valuation ring, $\mathfrak{m}_{K}$ for the maximal ideal, and $\tilde{K}=\mathcal{O}_{K}/\mathfrak{m}_{K}$ for the residue field. If $a\in\mathcal{O}_{K}$, we write $\tilde{a}$ for the reduction of $a$ — i.e., its image in the residue field.
Now let $L/K$ be a finite extension of $K$. Then $v$ extends uniquely to a valuation on $L$, and $L$ is complete and discretely valued with respect to the extended valuation. We will abuse notation by writing $v$ for this extension. The degree of the residue field extension is denoted by $\mathfrak{f}(L/K)$, and it satisfies $\mathfrak{f}(L/K)=\mathfrak{f}_{s}(L/K)\cdot\mathfrak{f}_{i}(L/K)$, where $\mathfrak{f}_{s}(L/K)$ and $\mathfrak{f}_{i}(L/K)$ are the separable and inseparable degrees of this extension, respectively. Write $\mathfrak{e}(L/K)$ for the ramification index $[v(L^{\times}):v(K^{\times})]$. As $K$ is complete, we have the following fundamental equality
$\mathfrak{e}(L/K)\cdot\mathfrak{f}_{i}(L/K)\cdot\mathfrak{f}_{s}(L/K)=[L:K].$
Definition 2.1.
If $L/K$ is a finite extension of complete discretely valued fields, then $L/K$ is
  • weakly unramified if $\mathfrak{e}(L/K)=1$;
  • unramified if $\mathfrak{f}_{i}(L/K)=\mathfrak{e}(L/K)=1$;
  • weakly totally ramified if $\mathfrak{f}_{s}(L/K)=1$; and
  • totally ramified if $\mathfrak{f}_{s}(L/K)=\mathfrak{f}_{i}(L/K)=1$.
If $L/K$ is weakly totally ramified and $\mathfrak{e}(L/K)$ is a power of the residue characteristic of $K$, then $L/K$ is weakly totally wildly ramified. (We use the convention that $10^{0}=$ 1 in order to include the case where $K$ has residue characteristic zero and $L=K$.)
Remark 2.2.
It is entirely standard that “unramified” means “trivial ramification index and separable residue extension”. Following [21, Section 0EXQ], a field is “weakly unramified” if it is unramified when we ignore separability issues in the residue field. We mimic this notion in the definitions of totally ramified and weakly totally ramified.
To address the discomfort that often accompanies working with inseparable extensions, we show that they behave as one might hope with respect to residue extensions and value groups, and that they may be approximated by separable extensions. The latter is well known [1], but we give a simple constructive argument for those interested in explicit calculation.
Proposition 2.3.
Let $K$ be a complete discretely valued field. Any finite purely inseparable extension of $K$ is weakly totally wildly ramified.
Proof.
Let $L/K$ be a purely inseparable extension. Then $[L:K]=\mathfrak{f}(L/K)\mathfrak{e}(L/K)$, so that all of the ramification is wild. Let $\tilde{L}/\tilde{K}$ be the residue extension. Suppose that $\mathfrak{f}_{s}(L/K)>1$. Then there is $\tilde{\alpha}\in\tilde{L}$ such that $\tilde{K}(\tilde{\alpha})/\tilde{K}$ is nontrivial and separable. Write $\tilde{g}$ for the minimal polynomial of $\tilde{\alpha}$, and let $g\in K[z]$ be an irreducible polynomial of the same degree whose coefficients reduce to that of $\tilde{g}$. By Hensel’s lemma, $g$ has a root $\alpha\in L$. But $g$ has no multiple root, else $\tilde{g}$ does. It follows that $K(\alpha)$ is a nontrivial separable subextension of $L/K$, a contradiction. ∎
In order to choose the extension $L/K$ to be separable in the statement of Ramified Approximation, we need to show that small perturbations of inseparable polynomials are separable.
Proposition 2.4.
If $f\in K[z]$ is an irreducible polynomial with a multiple root, then the roots of the separable polynomial $f(z)+cz$ converge uniformly to those of $f(z)$ as $v(c)\to\infty$. In particular, the separable closure $K_{\mathrm{s}}$ is dense in $K_{\mathrm{a}}$.
Proof.
We may assume $p>0$ is the characteristic of $K$. Since $f$ has a multiple root, we may write $f(z)=g(z^{q})$ with $q=p^{r}\geq p$ and some separable polynomial $g\in K[z]$. For $c\in K^{\times}$, set
$h_{c}(z)=f(z)+cz.$
Let $\alpha\in K_{\mathrm{a}}$ be a root of $f$. Write
$f(z)=a_{q}(z-\alpha)^{q}+a_{q+1}(z-\alpha)^{q+1}+\cdots,$
with $a_{i}\in K_{\mathrm{a}}$. Then
$h_{c}(z+\alpha)=c\alpha+cz+a_{q}z^{q}+a_{q+1}z^{q+1}+\cdots.$
If $v(c)$ is sufficiently large, then the Newton polygon for $h_{c}(z+\alpha)$ shows that $h_{c}$ has $q$ roots of the form $\alpha+u$ with
$v(u)=\frac{1}{q}\left(v(c\alpha)-v(a_{q})\right)\approx\frac{1}{q}v(c).\qed$
Remark 2.5.
Using MacLane’s method of approximants as described in §4, one can show that the fields $K[z]/(f)$ and $K[z]/(f+cz)$ have isomorphic residue fields and identical value groups when $v(c)$ is sufficiently large.
The following result was observed when $K$ is a finite extension of $\mathbb{Q}_{p}$ in [14, Prop. 1.2.4], though the proof is more general.
Lemma 2.6.
Let $L/K$ be an unramified extension of complete discretely valued fields. Suppose $D$ is a $G_{K}$-invariant disk containing an element of $L$. Then $D$ contains an element of $K$.
Proof.
Let $\pi$ be a uniformizer for $K$, and hence also for $L$. Let $S\subset\mathcal{O}_{L}$ be a system of coset representatives for the residue field of $L$ such that $0\in S$ and $s\in\mathcal{O}_{K}$ if and only if $\tilde{s}\in tildeK$, the residue field of $K$. Write $D=D(a,r)$ for some $a\in L$, and write
$a=a_{N}\pi^{N}+a_{N+1}\pi^{N+1}+\cdots,$
where the $a_{i}\in S$ for all $i$, and $N\in\mathbb{Z}$. If all $a_{i}\in K$, then $a\in K$. Otherwise, let $n$ be the minimum index such that $a_{i}\in K$ for all $i<n$ and $a_{n}\not\in K$. Set $b=\sum_{i<n}a_{i}\pi^{i}\in K$. Then $v(a-\sigma(a))\geq n$ for all $\sigma\in G_{K}$, with equality for any $\sigma$ that acts nontrivially on $a_{n}$. Now $n\geq r$ since $D$ is $G_{K}$-invariant, and it follows that
$v(b-a)=v(a_{n}\pi^{n}+\cdots)=n\geq r.$
That is, $b\in D\cap K$. ∎
3 The analytic affine line over $K$
Berkovich defined the analytic affine line in terms of multiplicative seminorms [4, §1.5]. In order to connect this theory with MacLane’s work, we need to recast the analytic affine line in the language of semivaluations. We also extend Berkovich’s classification of points to the non-algebraically closed setting.
For this paper, a $K$-semivaluation on $K[z]$ is a function $V\colon K[z]\to\mathbb{R}\cup\{\infty\}$ such that $V|_{K}=v$, and for all $f,g\in K[z]$, we have
$V(fg)=V(f)+V(g)\qquad\text{and}\qquad V(f+g)\geq\min(V(f),V(g)).$
(We will use capital letters to denote semivaluations on a polynomial ring, to distinguish from the valuation $v$ on the base field.) If, in addition, $V(f)=\infty\Rightarrow f=0$, then we will call $V$ a $K$-valuation.
For $f\in K[z]$ an irreducible polynomial, we define a $K$-semivaluation $V_{f,\infty}$ by
$V_{f,\infty}(g)=v\left(g(\alpha)\right),$
where $\alpha\in K_{\mathrm{a}}$ is a root of $f$. As $G_{K}$ acts on $K_{\mathrm{a}}$ by isometries, this definition is independent of the choice of $\alpha$.
Proposition 3.1.
Let $V$ be a $K$-semivaluation on $K[z]$, and suppose there exists a nonzero polynomial $g$ such that $V(g)=\infty$. Then $V=V_{f,\infty}$, where $f$ is a monic irreducible factor of $g$.
Proof.
The set $\mathfrak{p}_{V}=\{g\in K[z]\colon V(g)=\infty\}$ is a nonzero prime ideal of $K[z]$. Let $f$ be the monic irreducible generator of $\mathfrak{p}_{V}$. Then $V$ determines a valuation on the quotient field $L=K[z]/(f)$. As $V|_{K}=v$ and $K$ is complete, the valuation on $L$ is uniquely determined by $v$. But $V_{f,\infty}$ also determines a valuation on $L$, so by uniqueness, we have $V=V_{f,\infty}$. ∎
Consider $\mathbf{A}^{1,\mathrm{an}}_{K}$, the analytic affine line over $K$ in the sense of Berkovich [4, §1.5]. By definition, the points of $\mathbf{A}^{1,\mathrm{an}}_{K}$ are the multiplicative seminorms on $K[z]$ that extend the norm $\|\cdot\|:=\exp(-v(\cdot))$ on $K$. Taking $-\log(\cdot)$, we see that it is equivalent to describe the points of $\mathbf{A}^{1,\mathrm{an}}_{K}$ as the $K$-semivaluations on the polynomial ring $K[z]$. The topology is the weakest one such that every map $f\colon\mathbf{A}^{1,\mathrm{an}}_{K}\to\mathbb{R}\cup\{\infty\}$ given by $V\mapsto V(f)$ is continuous. (The connected neighborhoods of $\infty$ are of the form $(t,\infty]$.)
Given a $K$-semivaluation $V$ on $K[z]$, the set $\mathfrak{p}_{V}=\{f\in K[z]\colon V(f)=\infty\}$ is a prime ideal. The $K$-semivaluation $V$ defines a valuation on the domain $K[z]/\mathfrak{p}_{V}$, and hence also on its fraction field. Define $\widetilde{\mathcal{H}({V})}$ to be the associated residue class field. (This is Berkovich’s notation; in MacLane’s notation we have $\Lambda(V)=\widetilde{\mathcal{H}({V})}$.)
The space $\mathbf{A}^{1,\mathrm{an}}_{K}$ admits a partial order: we write $V\preceq W$ if and only if $V(f)\leq W(f)$ for all $f\in K[z]$. We will write $V\prec W$ to mean that $V\preceq W$ and $V\neq W$. Proposition 3.1 shows that the $K$-semivaluations $V_{f,\infty}$ are maximal for the partial order.
Remark 3.2.
Note that $V\prec W$ does not imply that $V(f)<W(f)$ for every polynomial $f$.
Recall $\mathbb{C}_{K}$ is the completion of an algebraic closure of $K$, and $G_{K}$ is the group of automorphisms of $\mathbb{C}_{K}$ that fix $K$. As $G_{K}$ acts by isometries for the unique extension of the valuation to $K_{\mathrm{a}}$, this allows us to extend $v$ to the completion $\mathbb{C}_{K}$ of $K_{\mathrm{a}}$. Since $v(K^{\times})=\mathbb{Z}$, we have $v(\mathbb{C}_{K}^{\times})=\mathbb{Q}$.
Any $\mathbb{C}_{K}$-semivaluation $\bar{V}$ on $\mathbb{C}_{K}[z]$ can be restricted to a $K$-semivaluation $K[z]$, from which we obtain a continuous map
$\operatorname{pr}\colon\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}\to\mathbf{A}^{1,\mathrm{an}}_{K}.$
The map $\operatorname{pr}$ preserves the partial order. The $G_{K}$-action extends to the affine line $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$, and we see that $\operatorname{pr}$ induces an isomorphism of locally ringed spaces
$\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}/G_{K}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\mathbf{A}^{1,\mathrm{an}}_{K}.$
(See [4, Cor. .3.6].) In particular, since $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ is path-connected [2, Cor. 1.14], so is $\mathbf{A}^{1,\mathrm{an}}_{K}$. One can even show that $\mathbf{A}^{1,\mathrm{an}}_{K}$ is uniquely path-connected by copying the proof for $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$.
Given a disk $D(a,r)$ inside $\mathbb{C}_{K}$, the association
$f\mapsto\inf_{b\in D(a,r)}v(f(b))$
defines a $\mathbb{C}_{K}$-semivaluation. (This is a consequence of the maximum principle in rigid geometry.) We write $\zeta_{a,r}$ for the point of $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ corresponding to the infimum semivaluation on $D(a,r)$. The associated semivaluation will be denoted $V_{a,r}$. (N.B. — Our notation $\zeta_{a,r}$ is nonstandard in that it is tailored to valuations.) Berkovich classified the points of $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ into four types [2, §2.1].
  • Type I: $\zeta_{a,\infty}$ for some $a\in\mathbb{C}_{K}$.
  • Type II: $\zeta_{a,r}$ for some $a\in\mathbb{C}_{K}$ and $r\in\mathbb{Q}$.
  • Type III: $\zeta_{a,r}$ for some $a\in\mathbb{C}_{K}$ and $r\in\mathbb{R}\smallsetminus\mathbb{Q}$.
  • Type IV: A limit of points $(\zeta_{a_{i},r_{i}})_{i\geq 0}$, where the associated sequence of disks $\left(D(a_{i},r_{i})\right)_{i\geq 0}$ is decreasing and has empty intersection.
A point $x$ is of type I if and only if the prime ideal $\mathfrak{p}_{x}$ is nonzero. A point $x$ is of type II if and only if the associated residue field $\widetilde{\mathcal{H}({x})}$ is of transcendence degree 1 over ~${\tilde{\mathbb{C}}}_{K}=(\tilde{K})_{\mathrm{a}}$.
The type of a point is preserved under the action of the automorphism group $G_{K}$, so we may extend Berkovich’s classification to $\mathbf{A}^{1,\mathrm{an}}_{K}$.
Definition 3.3.
A point $x\in\mathbf{A}^{1,\mathrm{an}}_{K}$ is of type $j$ (with $j=$ I, II, III, or IV) if each point of $\operatorname{pr}^{-1}(x)\subset\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ is of type $j$.
We want to give a more intrinsic description of the points of $\mathbf{A}^{1,\mathrm{an}}_{K}$ involving infimum semivaluations, as we have in the algebraically closed setting. To that end, for each irreducible $\phi\in K[z]$ and each $s\in\mathbb{R}\cup\{\infty\}$, define
$D(\phi,s)=\{b\in\mathbb{C}_{K}\colon\phi(b)\geq s\}.$
Following Rüth, we call these sets diskoids [17, §4.4.1]. (If $s\in\mathbb{Q}$ is finite, then $D(\phi,s)$ is a $K$-affinoid domain in the sense of rigid geometry.) If $s$ is sufficiently small, then $D(\phi,s)$ is a disk containing the roots of $\phi$. More generally, we now show that a diskoid is a $G_{K}$-orbit of disks.
Lemma 3.4.
Let $\phi\in K[z]$ be a nonconstant polynomial with roots $a_{1},\ldots,a_{n}$, listed with multiplicity. Let $a$ be any root of $\phi$. The function $M_{\phi,a}\colon\mathbb{R}\cup\{\infty\}\to\mathbb{R}\cup\{\infty\}$ given by
$r\mapsto\sum_{i=1}^{n}\min\{r,v(a_{i}-a)\}$
is continuous, piecewise affine, strictly increasing, and bijective.
Proof.
$M_{\phi,a}$ is a sum of functions that are continuous, piecewise affine, and non-decreasing, so it inherits all of those properties. Moreover, $a=a_{j}$ for some $j$, so the function $\min(r,v(a_{j}-a))=r$ is strictly increasing. It follows that $M$ is also strictly increasing. For $r$ sufficiently small, we see $M_{\phi,a}=r\deg(\phi)$. Also, $M_{\phi,a}(r)\to\infty$ as $r\to\infty$. Hence, $M_{\phi,a}$ is onto. ∎
If $\phi\in K[z]$ is irreducible, $M_{\phi,a}=M_{\phi,b}$ for any distinct roots $a,b$ of $\phi$. In that case, we will drop the root from the subscript and simply write $M_{\phi}$.
Proposition 3.5.
Let $\phi\in K[z]$ be irreducible and monic of degree $n$. Write $a_{1},\ldots,a_{n}$ for the roots of $\phi$, counted with multiplicity. Let $r,s\in\mathbb{R}\cup\{\infty\}$ be such that $M_{\phi}(r)=s$. Then
$D(\phi,s)=\bigcup_{i=1}^{n}D(a_{i},r).$
Proof.
If $r=s=\infty$, then the result is clear. In the remainder of the proof, we assume that $r,s\neq\infty$.
Write $\phi(z)=\prod(z-a_{i})$. Write $B=D(a_{1},r)$. If $b\in B$, then we have
$v(\phi(b))=\sum_{i=1}^{n}v(b-a_{i})\geq\sum_{\begin{subarray}{c}1\leq i\leq n\\ a_{i}\in B\end{subarray}}r+\sum_{\begin{subarray}{c}1\leq i\leq n\\ a_{i}\not\in B\end{subarray}}v(a_{1}-a_{i})=\sum_{i=1}^{n}\min\{r,v(a_{1}-a_{i})\}=s.$
That is, $b\in D(\phi,s)$. Since $D(\phi,s)$ is $G_{K}$-invariant, we find that $D(\phi,s)\supseteq\bigcup_{i}D(a_{i},r)$.
Now suppose that $b\not\in D(a_{i},r)$ for any $i$. If necessary, re-index the roots of $\phi$ so that
$v(b-a_{1})\geq v(b-a_{2})\geq\cdots\geq v(b-a_{n}).$
As before, set $B=D(a_{1},r)$. If $a_{i}\in B$, then $v(a_{i}-a_{1})\geq r$, and we find
$v(b-a_{i})<r=\min\{r,v(a_{i}-a_{1})\}.$
If $a_{i}\not\in B$, then $v(a_{i}-a_{1})<r$, so that
$v(b-a_{i})=\min\{v(b-a_{i}),v(b-a_{1})\}\leq v(a_{i}-a_{1})=\min\{r,v(a_{i}-a_{1})\}.$
Combining these observations yields
$v(\phi(b))=\sum_{i=1}^{n}v(b-a_{i})<\sum_{i=1}^{n}\min\{r,v(a_{i}-a_{1})\}=s.$
That is, $b\not\in D(\phi,s)$, and hence $D(\phi,s)\subseteq\bigcup_{i}D(a_{i},r)$. ∎
Proposition 3.6.
Let $\phi\in K[z]$ be irreducible, and let $s\in\mathbb{R}\cup\{\infty\}$. The map
$V_{\phi,s}(f):=\inf_{b\in D(\phi,s)}v(f(b))$
is a $K$-semivaluation on $K[z]$. Write $\zeta_{\phi,s}$ for the corresponding point of $\mathbf{A}^{1,\mathrm{an}}_{K}$. If
$D(\phi,s)=\bigcup_{i=1}^{n}D(a_{i},r)$
as in Proposition 3.5, then $\operatorname{pr}^{-1}(\zeta_{\phi,s})=\{\zeta_{a_{1},r},\ldots,\zeta_{a_{n},r}\}$.
Proof.
For $i=1,\ldots,n$, the $\mathbb{C}_{K}$-semivaluations $\zeta_{a_{i},r}$ agree when restricted to $K[z]$. Indeed, for $f\in K[z]$ and $\sigma\in G_{K}$, we have
$V_{a_{i},r}(f)$ $=\inf_{b\in D(a_{i},r)}v(f(b))$
$=\inf_{b\in D(a_{i},r)}v(\sigma(f(b)))$
$=\inf_{b\in D(a_{i},r)}v(f(\sigma(b)))$
$=\inf_{b\in\sigma(D(a_{i},r))}v(f(b))=V_{a_{j},r}(f),$
where $\sigma(a_{i})=a_{j}$. It follows that for $f\in K[z]$, we have
$\inf_{b\in D(\phi,s)}v(f(b))=\inf_{\begin{subarray}{c}1\leq i\leq n\\ b\in D(a_{i},r)\end{subarray}}v(f(b))=V_{a_{1},r}(f).$
In particular, $V_{\phi,s}$ is a $K$-semivaluation. We have also shown that $\{\zeta_{a_{1},r},\ldots,\zeta_{a_{n},r}\}\subset\operatorname{pr}^{-1}(\zeta_{\phi,s})$.
The opposite inclusion follows if we can show that $G_{K}$ acts transitively on the set of extensions of $V_{\phi,s}$ to $\mathbb{C}_{K}[z]$. Apply [11, Prop. II.9.1] to the infinite field extension $\mathbb{C}_{K}(z)/K(z)$, where $V_{\phi,s}$ is the valuation on $K(z)$. ∎
Proposition 3.6 allows us to give the desired classification of points of $\mathbf{A}^{1,\mathrm{an}}_{K}$:
  • Type I: $\zeta_{\phi,\infty}$ for some irreducible $\phi\in K[z]$.
  • Type II: $\zeta_{\phi,s}$ for some irreducible $\phi\in K[z]$ and $s\in\mathbb{Q}$.
  • Type III: $\zeta_{\phi,s}$ for some irreducible $\phi\in K[z]$ and $s\in\mathbb{R}\smallsetminus\mathbb{Q}$.
  • Type IV: A limit of points $(\zeta_{\phi_{i},s_{i}})_{i\geq 0}$, where the associated sequence of diskoids $\left(D(\phi_{i},s_{i})\right)_{i\geq 0}$ is decreasing and has empty intersection.
Remark 3.7.
If $\phi$ is irreducible and $c\in K^{\times}$, then $\zeta_{c\phi,s}=\zeta_{\phi,s-v(c)}$. In particular, nothing is lost in the above classification of points by assuming that the polynomial $\phi$ is monic.
4 MacLane’s method of approximants
Nearly a century ago, MacLane introduced an efficiently computable description of an arbitrary $K$-semivaluation on the polynomial ring $K[z]$ [8, 9]. We give a very brief introduction to set notation. MacLane does not require that $K$ be complete in his work, though our discussion is greatly simplified by doing so. See [9, §2] for full definitions and a quick-start guide, or see [8] for a more thorough treatment.
The main object of study in MacLane’s theory is the inductive semivaluation. A first stage inductive semivaluation $V_{1}$ arises from the standard semivaluation on a Tate algebra: there is $a\in K$ and $\mu_{1}\in\mathbb{R}\cup\{\infty\}$ such that
$V_{1}\left(\sum c_{m}(z-a)^{m}\right)=\min_{m}v(c_{m})+m\cdot\mu_{1}$
for each polynomial $\sum c_{m}(z-a)^{m}\in K[z]$. We write $V_{1}=[v,V_{1}(z-a)=\mu_{1}]$.
For $i>1$, an $i$th stage inductive semivaluation $V_{i}$ is defined as an augmentation of an $(i-1)$th stage valuation $V_{i-1}$: there is a (monic) key polynomial $\phi_{i}$ and a key value $\mu_{i}>V_{i-1}(\phi_{i})$ such that
$V_{i}\left(\sum c_{m}\phi_{i}^{m}\right)=\min_{m}V_{i-1}(c_{m})+m\cdot\mu_{i}.$
Here $c_{m}\in K[z]$ is a polynomial of degree strictly less than $\phi_{i}$; any polynomial $f\in K[z]$ admits a unique expansion of the form $\sum c_{m}\phi_{i}^{m}$. We write $V_{i}=[V_{i-1},V_{i}(\phi_{i})=\mu_{i}]$.
Given a $K$-semivaluation $W$ on $K[z]$, there exists a sequence of inductive valuations $V_{1}\prec V_{2}\prec V_{3}\prec\cdots$ such that $W=\lim V_{i}$ [8, Thm. 8.1]. In general, this sequence may be finite or infinite, and the description is not entirely constructive. However, if $f$ is an irreducible polynomial and $W=V_{f,\infty}$, then MacLane’s method of approximants [9] is constructive. We briefly recall the main steps:
  1. 1.
    Choose the first key polynomial to be $z-a$ for some $a\in K$. Write $f=\sum c_{j}(z-a)^{j}$ with $c_{j}\in K$. The lower convex hull of the points $(j,v(c_{j}))$ has a unique slope $m$; let the first key value be $\mu_{1}=-m$. Set $V_{1}=[v,V_{1}(z-a)=\mu_{1}]$.
  2. 2.
    Assume we have determined $V_{i}$, with key polynomial $\phi_{i}$ and key value $\mu_{i}$. If $\mu_{i}=\infty$, we are done. Otherwise, the residue field $\widetilde{\mathcal{H}({V_{i}})}$ is a rational function field $k_{i}(y)$, where $k_{i}$ is an extension of the residue field of $K$.
    • Find a polynomial $g$ with $\deg(g)<\deg(\phi_{i})$ such that $V_{i}(g)=-V_{i}(f)$. Let $\psi$ be an irreducible polynomial factor of the image of $gf$ in $\widetilde{\mathcal{H}({V_{i}})}$.
    • Let $\phi_{i+1}$ be a key polynomial for $V_{i}$ that is also a lift of $\psi$ in $K[z]$. (We may choose $\phi_{i+1}$ to be separable if we like since it is only well defined up to $V_{i}$-equivalence.)
    • Write $f=\sum c_{j}\phi_{i+1}^{j}$, where $c_{j}\in K[z]$ and $\deg(c_{j})<\deg(\phi_{i+1})$. The lower convex hull of the points $(j,V_{i}(c_{j}))$ has a unique slope $m$, and we take $\mu_{i+1}=-m$. Define $V_{i+1}=[V_{i},V_{i+1}(\phi_{i+1})=\mu_{i+1}]$.
The method of approximants constructs a particular sequence of points $\zeta_{1},\zeta_{2},\zeta_{3},\ldots\in\mathbf{A}^{1,\mathrm{an}}_{K}$ that converges to $\zeta_{f,\infty}$. Since $\mathbf{A}^{1,\mathrm{an}}_{K}$ is uniquely path-connected, each point $\zeta_{i}$ can be thought of as a blaze on the trail leading to $\zeta_{f,\infty}$. At each blaze, a new key polynomial $\phi$ is chosen such that $\zeta_{\phi,\infty}$ lies in the same direction as $\zeta_{f,\infty}$. One can show that the new key value $\mu$ is chosen so that $\zeta_{f,\infty}$ and $\zeta_{\phi,\infty}$ do not lie in the same direction at $\zeta_{\phi,\mu}$; that is, $\zeta_{\phi,\mu}$ is the maximal point $x\in\mathbf{A}^{1,\mathrm{an}}_{K}$ such that $x\preceq\zeta_{f,\infty}$ and $x\preceq\zeta_{\phi,\infty}$.
4.1 Diskoids and inductive valuations
Our primary goal in this section is to exhibit a correspondence between infimum semivaluations on diskoids and inductive semivaluations.
Proposition 4.1.
Let $V$ be an inductive semivaluation on $K[z]$, with key polynomial $\phi$ and key value $\mu$. Then for each $f\in K[z]$,
$V(f)=\inf_{b\in D(\phi,\mu)}v(f(b)).$
That is, $V$ corresponds to the point $\zeta_{\phi,\mu}\in\mathbf{A}^{1,\mathrm{an}}_{K}$.
Rüth proved this result in the case where $\mu$ is rational — i.e., $V$ corresponds to a type II point of $\mathbf{A}^{1,\mathrm{an}}_{K}$ [17, §4.4.2]. Using the path connectedness of $\mathbf{A}^{1,\mathrm{an}}_{K}$, we improve his argument and extend it to the case of irrational and infinite key values. This will require some preparation.
We begin with an easier result showing that a point of $\mathbf{A}^{1,\mathrm{an}}_{K}$ of type I, II, or III can be represented by some inductive semivaluation.
Proposition 4.2.
Let $\zeta_{f,s}$ be a point of $\mathbf{A}^{1,\mathrm{an}}_{K}$ of type I, II, or III with $f$ monic and irreducible. Then the corresponding semivaluation $V_{f,s}$ can be represented as an inductive semivaluation. If $s=\infty$, then we may take the final key polynomial to be $f$.
Proof.
We begin with the case of a type I point $\zeta_{f,\infty}$. If $f$ is linear, say $f=z-a$ for $a\in K$, then $V_{1}=[v,V(z-a)=\infty]=V_{f,\infty}$. If instead $\deg(f)>1$, then the method of approximants gives a sequence of inductive semivaluations $V_{1}\prec V_{2}\prec V_{3}\prec\cdots$ such that either $V_{f,\infty}=V_{n}$ for some $n\geq 2$, or else $V_{f,\infty}=\lim V_{i}$. We treat these two cases separately.
Suppose first that $V_{f,\infty}=V_{n}$ for some $n$. Then $V_{n}(\phi_{n})=\infty$, where $\phi_{n}$ is the $n$th key polynomial. This implies $f=\phi_{n}$ (Prop. 3.1).
Next suppose that we have an infinite sequence of inductive valuations such that $V_{f,\infty}=\lim V_{n}$. For $n$ sufficiently large, this means $V_{n}$ has the same value group as $K[z]/(f)$, and its residue field $\widetilde{\mathcal{H}({V_{n}})}$ is purely transcendental over the residue field of $K[z]/(f)$. It follows that $\deg(\phi_{n})=\deg(f)$. (If $f$ is separable, an estimate can be given for the size of $n$ [9, p.506].) We claim that $f$ is a key polynomial for $V_{n}$. Indeed, $f$ is monic, so $f=\phi+c$ for some $c\in K[z]$ with degree strictly smaller than $\deg(\phi)$. By construction, $\mu_{n}=v(c)$, so $V_{n}(f)=V_{n}(c)$. And $f$ is clearly equivalence irreducible since the image of $f/c$ in the residue ring is linear. MacLane shows that $f$ is a key polynomial for $V_{n}$ [8, Prop. 9.4]. We conclude that $W=[V_{n},W(f)=\infty]$ is a valid augmentation of $V_{n}$, and it equals $V_{f,\infty}$ by Proposition 3.1.
Now assume that $\zeta_{f,s}$ is a type II or III point of $\mathbf{A}^{1,\mathrm{an}}_{K}$. Let $t\in\mathbb{R}$ be the largest value such that the disk $D(z,t)$ contains $D(f,s)$. In particular, $\zeta_{z,t}\preceq\zeta_{f,s}\prec\zeta_{f,\infty}$. Take $V_{1}=[v,V_{1}(z)=t]$, and apply what we have already shown about type I points to represent $V_{f,\infty}$ as an inductive semivaluation $V_{n}$. This gives a sequence of inductive semivaluations $V_{1}\prec V_{2}\prec\cdots\prec V_{n}$ whose corresponding points $x_{1}<x_{2}<\cdots<x_{n}$ subdivide the segment $[\zeta_{z,t},\zeta_{f,\infty}]$.
If $V_{j}=V_{f,s}$ for some $j$, then we are finished. Suppose instead that $V_{j}\prec V_{f,s}\prec V_{j+1}$. Write $\phi=\phi_{j+1}$ for ease of notation. Since $\phi$ is a key polynomial for $V_{j}$, it follows that $W=[V_{j},W(\phi)=u]$ is a valid augmentation of $V_{j}$ for all $u>V_{j}(\phi)$. Moreover, $W$ traverses the interval $(V_{j},V_{j+1}]$ as $u$ increases along the real interval $(V_{j}(\phi),\mu_{j+1}]$. It follows that there is $u$ in this interval such that $W=V_{f,s}$, as desired. ∎
Lemma 4.3.
Let $V_{1}\prec\cdots\prec V_{n}$ be a sequence of augmented inductive semivaluations on $K[z]$. Write $\phi_{i}$ and $\mu_{i}$ for the key polynomial and key value of $V_{i}$, respectively. Then for each $i=2,\ldots,n$, we have
$\mu_{i}\deg(\phi_{i-1})>\mu_{i-1}\deg(\phi_{i}).$
Proof.
Write
$\phi_{i}=c_{m}\phi_{i-1}^{m}+\cdots+c_{1}\phi_{i-1}+c_{0},$
where $\deg(c_{j})<\deg(\phi_{i-1})$ for each $j$. Since $\phi_{i}$ is a key polynomial for $V_{i-1}$, we have $c_{m}=1$ and $V_{i-1}(\phi_{i})=V_{i-1}(\phi_{i-1}^{m})$ [8, Prop. 9.4]. It follows that $m\deg(\phi_{i-1})=\deg(\phi_{i})$, and
$\mu_{i}=V_{i}(\phi_{i})>V_{i-1}(\phi_{i})=m\mu_{i-1}.\qed$
The next result is due to Rüth [17, Lem. 4.53], though we fill a small gap in the proof with Lemma 4.3.
Lemma 4.4.
Let $V_{1},V_{2},\ldots,V_{n}$ be a sequence of augmented inductive semivaluations on $K[z]$. Write $\phi_{i}$ and $\mu_{i}$ for the key polynomial and key value of $V_{i}$, respectively. Then
$D(\phi_{n},\mu_{n})\subsetneq\cdots\subsetneq D(\phi_{2},\mu_{2})\subsetneq D(\phi_{1},\mu_{1}).$
Proof.
Write $D_{i}=D(\phi_{i},\mu_{i})$ for each $i$. Fix an index $i\geq 1$, and let us show that $D_{i+1}\subsetneq D_{i}$. For ease of notation, set $\phi=\phi_{i+1}$. We claim that the roots of $\phi$ lie inside $D_{i}$. Define $W=[V_{i},W(\phi)=\infty]$. Then if $\alpha$ is a root of $\phi$, we have
$\mu_{i}=V_{i}(\phi_{i})=W(\phi_{i})=v(\phi_{i}(\alpha)),$
so that $\alpha\in D_{i}$. Since $D_{i}$ is $G_{K}$-invariant, we conclude that all roots of $\phi$ lie in $D_{i}$.
Since $D(\phi,\infty)\subset D(\phi_{i},\mu_{i})$, we can write $D(\phi,s)=D(\phi_{i},\mu_{i})$ for some $s\in\mathbb{R}$. Write
$D(\phi,s)=B_{1}\cup\cdots\cup B_{d},$
a union of $G_{K}$-translates of a disk of radius $r=M_{\phi}^{-1}(s)$ (Prop. 3.5). Set $r_{1}=r$ and $r_{j}=v(B_{1}-B_{j})$ for $j=2,\ldots,d$. (As $B_{1}$ and $B_{j}$ are disjoint, the difference between any pair of elements from $B_{1}$ and $B_{j}$ has the same valuation.) Each disk must contain $\deg(\phi)/d$ roots of $\phi$, counted with multiplicity. If we write $a_{i}$ for a root of $\phi$, then we have
$s$ $=\sum_{i=1}^{\deg(\phi)}\min\{r,v(a_{1}-a_{i})\}$
$=\sum_{j=1}^{d}\ \sum_{\begin{subarray}{c}1\leq i\leq\deg(\phi)\\ a_{i}\in B_{j}\end{subarray}}\min\{r,v(a_{1}-a_{i})\}$
$=\frac{\deg(\phi)}{d}(r+r_{2}+\cdots+r_{d}).$
Applying this same argument to the decomposition $D(\phi_{i},\mu_{i})=B_{1}\cup\cdots\cup B_{d}$ shows that
$\mu_{i}=\frac{\deg(\phi_{i})}{d}(r+r_{2}+\cdots+r_{d}).$
Combining the last two displayed equations, we find that
$s=\mu_{i}\frac{\deg(\phi)}{\deg(\phi_{i})}.$
Then $s<\mu_{i+1}$ (Lem. 4.3), which means $D(\phi_{i},\mu_{i})=D(\phi,s)\supsetneq D(\phi,\mu_{i+1})$. ∎
Proof of Proposition 4.1.
Let $V_{1}\prec\cdots\prec V_{n}$ be a sequence of augmented inductive semivaluations on $K[z]$, where $V=V_{n}$. Write $\phi,\mu$ for the key polynomial and key value for $V_{n}$, respectively. We want to show that $V_{n}=V_{\phi,\mu}$.
To begin, we claim that $V_{\phi,\mu}\succeq V_{n}$. That is, if $b\in D(\phi,\mu)$, then $v(f(b))\geq V_{n}(f)$ for all $f\in K[z]$. Suppose first that $n=1$. Write $V_{1}=[v,V_{1}(z-a)\geq\mu]$ for some $a\in K$ and $\mu\in\mathbb{R}\cup\{\infty\}$. For $f\in K[z]$, we write $f=\sum c_{i}(z-a)^{i}$ with $c_{i}\in K$. For $b\in D(z-a,\mu)$, we have
$v(f(b))\geq\min_{i}\{v(c_{i})+i\cdot v(b-a)\}\geq\min_{i}\{v(c_{i})+i\mu\}=V_{1}(f).$
Now suppose the claim holds for all indices $i<n$. Write $V_{n}=[V_{n-1},V_{n}(\phi)=\mu]$. For $f\in K[z]$, we write $f=\sum c_{i}\phi^{i}$, where $c_{i}\in K[z]$ satisfies $\deg(c_{i})<\deg(\phi)$. For any $b\in D(\phi,\mu)$, we find that
$v(c_{i}(b))\geq V_{\phi_{n-1},\mu_{n-1}}(c_{i})\geq V_{n-1}(c_{i}),$
where the first inequality is a consequence of $D(\phi_{n-1},\mu_{n-1})\supsetneq D(\phi,\mu)$ (Lem. 4.4), and the second is our inductive hypothesis. The ultrametric inequality gives
$v(f(b))\geq\min_{i}\{v(c_{i}(b))+i\cdot v(\phi(b))\}\geq\min_{i}\{V_{n-1}(c_{i})+i\mu\}=V_{n}(f).$
Since $b\in D(\phi,\mu)$ is arbitrary, we have $V_{\phi,\mu}\succeq V_{n}$
Since $V_{n}\preceq V_{\phi,\mu}$, there is some $s\leq\mu$ with $V_{n}=V_{\phi,s}$. To complete the proof, it suffices to show that $V_{\phi,s}(\phi)=s$ for all $s\in\mathbb{R}$.
Evidently, the definition of the infimum semivaluation yields $V_{\phi,s}(\phi)\geq s$. Define $r=M_{\phi}^{-1}(s)$, and write $\phi(z)=\prod(z-a_{i})$. Proposition 3.5 gives $D(\phi,s)=\bigcup_{i}D(a_{i},r)$. If $r\in\mathbb{Q}=v(K_{\mathrm{a}}^{\times})$, then there exists $b\in D(a_{1},r)$ such that $v(b-a_{j})=r$ for each $a_{j}\in D(a_{1},r)$. This gives
$v(\phi(b))=\sum_{i=1}^{\deg(\phi)}v(b-a_{i})=\sum_{i=1}^{\deg(\phi)}\min\{r,v(a_{1}-a_{i})\}=s.$
That is, $V_{\phi,s}(\phi)=s$. If instead $r\not\in\mathbb{Q}$, then we may approximate $r$ by rational numbers and apply the preceding argument to produce $b\in D(a_{1},r)$ such that $v(\phi(b))\leq s+\varepsilon$ for any $\varepsilon>0$. So $V_{\phi,s}(\phi)=s$ in that case as well. The proof is complete. ∎
4.2 Local geometry at a type II point
Convention. In this section, we write $k$ for the residue field of $K$ (instead of $\tilde{K}$). Then the residue field of $\mathbb{C}_{K}$ is $k_{\mathrm{a}}$.
To motivate our need for local geometric considerations, we sketch the idea of the proof of Ramified Approximation. Let $f\in K[z]$ be an irreducible polynomial. Apply MacLane’s method of approximants to obtain a sequence of inductive semivaluations $V_{1}\prec V_{2}\prec\cdots\prec V_{n}=V_{f,\infty}$. This corresponds to a sequence of points $\zeta_{1}\prec\zeta_{2}\prec\cdots\prec\zeta_{n}=\zeta_{f,\infty}$ of $\mathbf{A}^{1,\mathrm{an}}_{K}$, which in turn corresponds to a descending sequence of diskoids $D_{1}\supsetneq D_{2}\supsetneq\cdots\supsetneq D_{n}$, each of which contains the roots of $f$. The minimal disk $D$ containing the roots of $f$ sits somewhere in this chain, say $D_{j}\supseteq D\supsetneq D_{j+1}$. In particular, for $1\leq i\leq j$, the diskoid $D_{i}$ is actually a disk, while for $i\geq j+1$, it is a nontrivial $G_{K}$-orbit of disks. This bifurcation can be detected in the way that “tangent vectors” collapse under the map $\operatorname{pr}\colon\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}\to\mathbf{A}^{1,\mathrm{an}}_{K}$.
The notion of tangent space at a point of $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ has been explored thoroughly; for a quick summary, see [6, §2.2.5]. We now extend these ideas to $\mathbf{A}^{1,\mathrm{an}}_{K}$.
Proposition 4.5.
Let $\zeta$ be a type II point of $\mathbf{A}^{1,\mathrm{an}}_{K}$, represented by an inductive valuation $V$. Write ${\widetilde{\mathcal{H}({V})}\cong\ell(x)}$ for the residue field of $V$, where $\ell$ is a finite extension of $k$, the residue field of $K$.
  • The places of $\ell(x)$ parameterize the connected components of $\mathbf{A}^{1,\mathrm{an}}_{K}\smallsetminus\{\zeta\}$.
  • The finite places of $\ell(x)$ parameterize equivalence classes of key polynomials for $V$.
Proof.
The infinite place of $\ell(x)$ corresponds to all valuations $W$ with $W\not\succeq V$. Each other place of $\ell(x)$ is described by the order of vanishing at an irreducible monic polynomial $\psi\in\ell[x]$. Then $\psi$ corresponds to a particular equivalence class of key polynomials $\phi$ for $V$ [8, Thm. 13.1] [^1] The equivalence class corresponding to $\psi=x$ was excluded from [8, Thm. 13.1] in order to use [8, Thm. 9.4], but loc.cit. contains an error. It does not allow for $\phi_{k}$ to be a key polynomial for $V_{k}$. See [9, Lem. 4.3]., and equivalent key polynomials give the same germ of an augmentation. (Two augmentations $W,W^{\prime}$ yield the same germ if there is an augmentation $W^{\prime\prime}$ of $V$ such that $W^{\prime\prime}\preceq W$ and $W^{\prime\prime}\preceq W^{\prime}$.) Finally, a germ at $V$ describes an equivalence class of augmentations $W$ with $V\prec W$, and hence a connected component of $\mathbf{A}^{1,\mathrm{an}}_{K}\setminus\{\zeta\}$. ∎
Let $\zeta\in\mathbf{A}^{1,\mathrm{an}}_{K}$ be a point. Write $T_{\zeta}$ for the set of connected components of $\mathbf{A}^{1,\mathrm{an}}_{K}\smallsetminus\{\zeta\}$; an element $\vec{w}\in T_{\zeta}$ will be called a tangent vector at $\zeta$. If we wish to emphasize that it is a subset of $\mathbf{A}^{1,\mathrm{an}}_{K}$, we will write $U_{\vec{w}}$.
Take $\alpha\in\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ to be a point, and let $\zeta=\operatorname{pr}(\alpha)\in\mathbf{A}^{1,\mathrm{an}}_{K}$ be its image. A tangent vector $\vec{v}\in T_{\alpha}$ corresponds to a connected component $U_{\vec{v}}$, and one sees that $\operatorname{pr}(U_{\vec{v}})$ is a connected component of $\mathbf{A}^{1,\mathrm{an}}_{K}\smallsetminus\{\zeta\}$. It follows that $\operatorname{pr}$ induces a well defined map $\operatorname{pr}_{*}\colon T_{\alpha}\to T_{\zeta}$. Given $\vec{w}\in T_{\zeta}$, we define the branch multiplicity $m(\zeta,\vec{w})$ to be the number of distinct vectors $\vec{v}\in T_{\alpha}$ such that $\operatorname{pr}_{*}(\vec{v})=\vec{w}$. As the notation suggests, this number is independent of the choice of $\alpha\in\operatorname{pr}^{-1}(\zeta)$ because $G_{K}$ acts transitively on this set and carries tangent vectors to tangent vectors. See Figure 1.
$\mathbf{A}^{1,\mathrm{an}}_{K}$$\zeta$$\vec{w}$$\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$$\alpha$$\operatorname{pr}$
Figure 1: An example showing the projection map locally collapsing $m(\zeta,\vec{w})=3$ branches at $\alpha\in\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ to a single branch at $\zeta\in\mathbf{A}^{1,\mathrm{an}}_{K}$. Here $\operatorname{pr}^{-1}(\zeta)$ consists of two points.
Keep the setup of the last paragraph. The map $\operatorname{pr}\colon\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}\to\mathbf{A}^{1,\mathrm{an}}_{K}$ is a morphism of locally ringed spaces, so it induces a $k$-algebra homomorphism of residue fields $\widetilde{\mathcal{H}({\zeta})}\to\widetilde{\mathcal{H}({\alpha})}$. If $\alpha$ and $\zeta$ are type II points, then in suitable coordinates this is a homomorphism $\ell(x)\to k_{{\mathrm{a}}}(y)$. To describe it precisely, we need some additional terminology.
Write $W$ for the valuation corresponding to $\zeta$; we may assume that $W$ is an inductive valuation (Prop. 4.2). If $W$ is a first stage inductive valuation, write $\Gamma=\mathbb{Z}$. Otherwise, $W$ is an augmentation of another inductive valuation $V$, and we write $\Gamma$ for the value group of $V$. If $f\in K[z]$ is a polynomial such that $W(f)\in\Gamma$, then there is an equivalence unit $g\in K[z]$ such that $W(fg)=0$ and $gg^{\prime}\equiv 1\pmod{V}$ for some $g^{\prime}\in K[z]$ [8, Lem. 9.1/3]. (Intuitively, equivalence units are the polynomials that have no root in the direction at $V$ determined by $W$.) A residual polynomial for $f$ is the image of $fg$ in $\widetilde{\mathcal{H}({W})}$. It is well defined up to multiplication by an element of the constant field of $\widetilde{\mathcal{H}({W})}$.
Let $\phi$ and $s$ be the key polynomial and key value for $W$, respectively. Write $\tau\geq 1$ for the relative ramification index of $W/V$; i.e., $\tau$ is the minimal positive integer such that $\tau s\in\Gamma$. Then $W(\phi^{\tau})\in\Gamma$, and we take an equivalence unit $g$ such that $W(g\phi^{\tau})=0$. Define $x$ to be the image of $g\phi^{\tau}$ in $\widetilde{\mathcal{H}({W})}$; it is our distinguished residual polynomial for $\phi^{\tau}$. Then $\widetilde{\mathcal{H}({W})}\cong\ell(x)$, where $\ell=k_{\mathrm{a}}\cap\widetilde{\mathcal{H}({W})}$ [8, Cor. 12.2].
Let $D(a,r)$ be a disk in the decomposition of $D(\phi,s)$ (Prop. 3.5), so that $\operatorname{pr}(V_{a,r})=V_{\phi,s}=W$. If $c\in K_{\mathrm{a}}$ is such that $v(c)=-r$, then $V_{a,r}(c(z-a))=0$, and we define $y$ to be the image of $c(z-a)$ in $\widetilde{\mathcal{H}({V_{a,r}})}$. If $D(a,r)$ contains exactly $d$ zeros of $\phi$, when counted with multiplicity, then the image of $g\phi^{\tau}$ in $\widetilde{\mathcal{H}({V_{a,r}})}$ is $\eta^{\tau}$ for some polynomial $\eta$ of degree $d$. We have thus proved that there is a choice of coordinates such that the composition
$\ell(x)\cong\widetilde{\mathcal{H}({V_{\phi,s}})}\longrightarrow\widetilde{\mathcal{H}({V_{a,r}})}\cong k_{\mathrm{a}}(y)$
is given by $x\mapsto\eta^{\tau}$, where $\eta\in k_{\mathrm{a}}[y]$ is a particular polynomial whose degree is the number of zeros of $\phi$ inside $D(a,r)$, counted with multiplicity.
Now let $\psi\neq x$ be a monic irreducible polynomial, which corresponds to a vector $\vec{w}\in T_{\zeta}$ (Prop. 4.5). Then pushing $\psi$ through the above map gives $\psi(\eta^{\tau})$. The distinct roots of this polynomial correspond to places of $k_{\mathrm{a}}(y)$, which in turn correspond to tangent vectors at $\zeta_{a,r}$ that map to $\vec{w}$. Counting these roots, we have proved that
$m(\zeta,\vec{w})=\deg_{\mathrm{s}}(\psi)\cdot\deg_{\mathrm{s}}(\eta)\cdot\tau^{(p)},$ (4.1)
where $\deg_{\mathrm{s}}$ denotes the separable degree of a polynomial, and $\tau^{(p)}=\tau/p^{\operatorname{ord}_{p}(\tau)}$ is the $p$-free part of $\tau$. We record an important consequence of this formula.
Proposition 4.6.
Let $V_{n}$ be an inductive valuation with key polynomial $\phi_{n}$ and key value $\mu_{n}\in\mathbb{Q}$. Let $\tau$ be the relative ramification index of $V_{n}$ over $V_{n-1}$. Let $\phi\in K[z]$ be a key polynomial for $V_{n}$ such that $\phi\not\sim\phi_{n}\pmod{V_{n}}$. For each $\mu>V_{n}(\phi)$, we have $|\operatorname{pr}^{-1}(\zeta_{\phi,\mu})|>|\operatorname{pr}^{-1}(\zeta_{\phi_{n},\mu_{n}})|$ if one of the following holds:
  1. 1.
    A residual polynomial for $\phi\pmod{V_{n}}$ has at least two distinct roots in $k_{\mathrm{a}}$, or
  2. 2.
    $\tau>1$ and $\tau$ is not a power of the residue characteristic of $K$.
Proof.
Write $\zeta=\zeta_{\phi_{n},\mu_{n}}$ and let $\alpha\in\operatorname{pr}^{-1}(\zeta)$. Let $\vec{w}\in T_{\zeta}$ be the tangent vector corresponding to the equivalence class of key polynomials containing $\phi$. We claim that the branch multiplicity $m(\zeta,\vec{w})>1$. Indeed, let $\psi$ be a residual polynomial for $\phi$. Then condition 1 implies $\deg_{\mathrm{s}}(\psi)>1$, while condition 2 implies $\tau^{(p)}>1$; the claim follows from (4.1).
Now let $\zeta^{\prime}=\zeta_{\phi,\mu}$ for some $\mu>\mu_{n}$. Then $\zeta^{\prime}\in U_{\vec{w}}$. For each $\vec{v}\in T_{\alpha}$ such that $\operatorname{pr}_{*}(\vec{v})=\vec{w}$, we have $\operatorname{pr}(U_{\vec{v}})=U_{\vec{w}}$. It follows that
$\big{|}\operatorname{pr}^{-1}(\zeta^{\prime})\big{|}\geq m(\zeta,\vec{w})\ \big{|}\operatorname{pr}^{-1}(\zeta)\big{|}>\big{|}\operatorname{pr}^{-1}(\zeta)\big{|}.\qed$
5 Ramified Approximation
Convention. In this section, we write $k$ for the residue field of $K$ (instead of $\tilde{K}$). Then the residue field of $\mathbb{C}_{K}$ is $k_{\mathrm{a}}$.
We dispense with the (easier) case where $K$ has residue characteristic zero first, and then we spend the remainder of the section on the case of positive residue characteristic.
Proposition 5.1.
Suppose the residue characteristic of $K$ is zero. Let
$f=c_{n}z^{n}+c_{n-1}z^{n-1}+\cdots+c_{0}$
be an irreducible polynomial in $K[z]$, and let $B\subset\mathbb{C}_{K}$ be a disk containing the roots of $f$. Then $B$ contains the $K$-rational element $-c_{n-1}/(nc_{n})$.
Proof.
If $a=a_{1},\ldots,a_{n}$ are the distinct roots of $f$ in $K_{\mathrm{a}}$, then
$f=c_{n}\prod(z-a_{i})=c_{n}z^{n}-c_{n}(\sum a_{i})z^{n-1}+\cdots,$
and $-c_{n-1}/(nc_{n})=\frac{1}{n}\sum a_{i}\in K$. Write $b$ for this element. Let $r\in\mathbb{R}$ be such that $B=D(a,r)$. Then
$v(b-a)=v\left(\frac{1}{n}\sum_{i=1}^{n}a_{i}-a\right)=v\left(\sum_{i=1}^{n}(a_{i}-a)\right)\geq\min_{i}v(a_{i}-a)\geq r.$
That is, $b\in B$. ∎
Now we apply the method of approximants to locate weakly totally wildly ramified elements inside $G_{K}$-invariant disks.
Theorem 5.2.
Suppose the residue characteristic of $K$ is $p>0$. Let $f\in K[z]$ be irreducible, and let $B\subset\mathbb{C}_{K}$ be a disk containing the roots of $f$. Let $V_{1}\prec V_{2}\prec\cdots\prec V_{n}=V_{f,\infty}$ be approximants to the $K$-semivaluation $V_{f,\infty}$, with key polynomials $\phi_{1},\phi_{2},\ldots,\phi_{n}=f$. Take $j\leq n$ to be the maximum index such that
  1. 1.
    $\widetilde{\mathcal{H}({V_{j}})}\cap k_{\mathrm{a}}$ is purely inseparable, and
  2. 2.
    the ramification index of $V_{i}/v$ is a power of $p$ for all $i<j$.
Then the roots of the key polynomial $\phi_{j}$ are contained in $B$, and $K[z]/(\phi_{j})$ is weakly totally wildly ramified.
Remark 5.3.
The same approach applies when the residue characteristic of $K$ is zero, though now we simply look for the largest index $j$ such that the key polynomial for $V_{j}$ is linear. If $\phi_{j}=z-b$, then $b\in B$ is the element we seek.
Proof of Theorem 5.2.
Observe that $\widetilde{\mathcal{H}({V_{n}})}$ is the residue field of the extension $K[z]/(f)$. If a root of $f$ generates a weakly totally wildly ramified extension of $K$, then $\widetilde{\mathcal{H}({V_{n}})}$ is purely inseparable over $k$, and the result follows. This includes the case where $f$ is purely inseparable (Prop. 2.3), so we may assume in the remainder of the proof that $f$ is not purely inseparable. In particular, $B$ has finite radius.
The maximum index $j$ such that conditions 1 and 2 hold must satisfy $j<n$. For ease of notation, write $\phi=\phi_{j}$ and $\mu=\mu_{j}$, so that $V_{j}=[V_{j-1},V_{j}(\phi)=\mu]$. To see that $K[z]/(\phi)$ is weakly totally wildly ramified, consider the sequence of augmentations
$V_{1}\prec V_{2}\prec\cdots\prec V_{j-1}\prec W,$
where $W=[V_{j-1},W(\phi)=\infty]$. Then $\widetilde{\mathcal{H}({V_{j}})}\cap k_{\mathrm{a}}=\widetilde{\mathcal{H}({W})}\cap k_{\mathrm{a}}$, and this extension depends only on the residue ring for $V_{j-1}$ and the key polynomial $\phi$ [8, Thm. 12.1]. In particular, since $\widetilde{\mathcal{H}({W})}$ is the residue field of $K[z]/(\phi)$, we see that it is purely inseparable. The ramification index of $V_{j-1}/v$ is a power of $p$, and it agrees with the ramification index of $K[z]/(\phi)$ over $K$.
It remains to show that the roots of $\phi$ lie inside the disk $B$. Define $D^{\prime}=D(\phi,\mu)$. If $D^{\prime}\subset B$, then $B$ contains the roots of $\phi$ (Lem. 4.4), and we are finished. So suppose $D^{\prime}\not\subset B$. Since $D^{\prime}$ and $D(\phi_{1},\mu_{1})$ both contain the roots of $f$ (Lem. 4.4), it follows that
$B\subsetneq D^{\prime}\subseteq D(\phi_{1},\mu_{1}),$
as we can write all three of these disks as $D(f,s)$ for various choices of $s$. Since $D^{\prime}$ contains $B$, we see that $D(\phi_{i},\mu_{i})$ is a $G_{K}$-invariant disk for each $i\leq j$. Now we use the maximality of the index $j$.
Suppose first that condition 1 fails for $j~{}+~{}1$. Then the residue field of $V_{j+1}$ has nontrivial separable degree over $k$, and $|\operatorname{pr}^{-1}(V_{\phi_{j+1},t})|>|\operatorname{pr}^{-1}(V_{\phi,\mu})|$ for all $t>V_{j}(\phi_{j+1})$ (Prop. 4.6). As $V_{j+1}$ and the infimum valuation on $B$ lie in the same direction from $V_{j}$, this would imply $B$ is a nontrivial Galois orbit of disks. But $B$ is a single disk, so this is a contradiction.
Now suppose that condition 2 fails for $j+1$, which means the relative ramification index of $V_{j}/V_{j-1}$ is not a power of $p$. We obtain a contradiction just as in the previous paragraph.
Since either condition 1 or 2 must fail for the index $j+1$, we are forced to conclude that $D^{\prime}\subseteq B$, and that the roots of $\phi$ are contained in $B$. ∎
Example 5.4.
We caution the reader that, with the setup of Theorem 5.2, the diskoid corresponding to the approximant $V_{j}$ may lie strictly inside the minimal disk about the roots of $f$. For example, take $K=\mathbb{Q}_{2}$ and consider the polynomial
$f(z)=z^{4}+20z^{2}+292.$
The method of approximants gives $V_{1}=[\operatorname{ord}_{2},V_{1}(z)=1/2]$, $V_{2}=[V_{1},V_{2}(z^{2}+2)=4]$, and $V_{3}=V_{f,\infty}$. Theorem 5.2 tells us that $\sqrt{-2}$ lies inside $B$, the minimal disk about the roots of $f$. This minimal disk is $B=D(z^{2}+2,3)$, while the diskoid for $V_{2}$ is $D(z^{2}+2,4)$. See Figure 2.
$\zeta_{1}$$\zeta_{B}$$\zeta_{2}$$\zeta_{3}=\zeta_{f,\infty}$$\mathbf{A}^{1,\mathrm{an}}_{\mathbb{Q}_{2}}$$\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{2}}$$\operatorname{pr}$
Figure 2: A diagram of the positions of approximants for $f(z)=z^{4}+20z^{2}+292$ in $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{Q}_{2}}$ and their pre-images in $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{2}}$. Here $\zeta_{i}$ corresponds to the approximant $V_{i}$, while $\zeta_{B}$ corresponds to the infimum valuation on $B$. The minimum disk about the roots of $f$ is $D(z^{2}+2,3)$; the corresponding point downstairs, $\zeta_{B}$ is not an approximant. See Example 5.4
Proof of Ramified Approximation.
Suppose first that $K$ has residue characteristic zero. The result follows immediately from Proposition 5.1.
Now suppose that $K$ has residue characteristic $p>0$. Taking $B=D$, Theorem 5.2 gives a polynomial $\phi=\phi_{j}$ whose roots all lie in $D$. Let $\alpha\in K_{\mathrm{a}}$ be a root of $\phi$. Then $K(\alpha)/K$ is weakly totally wildly ramified. The construction of the approximants $V_{1}\prec V_{2}\prec\cdots\prec V_{n}=V_{f,\infty}$ gives a sequence of residue extensions $k\subset k_{1}\subset k_{2}\subset\cdots\subset\ell$, where the residue field of $K(\alpha)$ is $k_{j}$ [8, Thm. 12.1]. This proves the claim about residue fields. A similar statement holds for value group extensions [9, §6].
Finally, we note that in choosing key polynomials, we always have the flexibility to vary within an equivalence class. In particular, by adding an appropriate linear term $cz$ with $v(c)$ sufficiently large, we may assume that $\phi$ is separable (Prop. 2.4). That is, we may arrange for the extension $K(\alpha)/K$ to be separable. ∎
The Approximation Theorem, as stated, is sufficient to deduce Theorem B. In order to obtain some of the degree bounds in Theorem A, we need to look more closely at certain quartic polynomials. If $f\in K[z]$ is a separable polynomial, we say that the roots of $f$ are equispaced if $v(\alpha-\beta)$ is the same for any pair of distinct roots $\alpha,\beta$ of $f$.
Lemma 5.5.
Suppose $K$ has residue characteristic 2, and let $f\in K[z]$ be a separable quartic polynomial with equispaced roots. Then either $f$ admits a weakly totally ramified root, or else there is $\alpha\in K_{\mathrm{a}}$ such that
  • $K(\alpha)/K$ is separable and weakly totally ramified of degree dividing 2;
  • $\alpha$ lies in the minimal disk in $\mathbb{C}_{K}$ containing the roots of $f$; and
  • $v(\beta-\gamma)\in v(K(\alpha)^{\times})$ for every distinct pair of roots $\beta,\gamma$ of $f$.
Proof.
Assume that $f$ has no weakly totally ramified root. Without loss of generality, we may assume that $f$ is monic. Then one of the following must be true:
  1. 1.
    $f=f_{1}f_{2}$ for quadratic polynomials $f_{1},f_{2}\in K[x]$ and $K[z]/(f_{i})$ is unramified for $i=1,2$;
  2. 2.
    $f$ is irreducible and $K[z]/(f)$ is unramified;
  3. 3.
    $f$ is irreducible and $K[z]/(f)$ has ramification index 2 and separable residue degree $2$; or
  4. 4.
    $f$ is irreducible and $K[z]/(f)$ has separable residue degree 2 and inseparable residue degree 2.
Let $D$ be the minimal disk about the roots of $f$. By Ramified Approximation, we may choose $\alpha\in D$ such that $K(\alpha)/K$ is separable and weakly totally ramified of degree dividing 2. All that remains to show is that the difference between any pair of roots has rational valuation over $K(\alpha)$.
Suppose first that we are in case 1 or 2. Then a root of $f$ generates an unramified extension, and so we may take $\alpha\in K$ (Lem. 2.6). Define $g(z)=f(z+\alpha)$. Since the roots of $f$ are equispaced, the roots of $g$ must all have the same valuation $s$. If $s\not\in\mathbb{Z}$, then a root of $g$ generates a ramified extension of $K$, a contradiction.
Now suppose that we are in case 3 or 4. Let $B$ be the minimal disk containing the roots of $f$. Write $\zeta\in\mathbf{A}^{1,\mathrm{an}}_{K}$ for the point corresponding to the infimum valuation on $B$, and let $\xi\in\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ be the unique point such that $\operatorname{pr}(\xi)=\zeta$. Let $\vec{w}\in T_{\zeta}$ be the tangent vector such that $U_{\vec{w}}$ contains $\zeta_{f,\infty}$. Since the roots of $f$ are equispaced, there are precisely four vectors $\vec{v}_{i}\in T_{\xi}$ such that $\operatorname{pr}_{*}(\vec{v}_{i})=\vec{w}$.
Let $V_{1}\prec V_{2}\prec\cdots\prec V_{n}=V_{f,\infty}$ be a minimal representation of $V_{f,\infty}$ as an inductive semivaluation (Prop. 4.2). The key polynomials of a minimal representation have strictly increasing degrees [8, Lem. 15.1], so $n=2$ or $n=3$. Since the separable degree of the residue extension of $K[z]/(f)$ is nontrivial, there is an index $j$ such that a residual polynomial for the key polynomial $\phi_{j+1}$ has nontrivial separable degree. If $\zeta_{j}\in\mathbf{A}^{1,\mathrm{an}}_{K}$ corresponds to $V_{j}$ and $\vec{u}\in T_{\zeta_{j}}$ corresponds to (the equivalence class of) $\phi_{j+1}$, the branch multiplicity formula (4.1) shows that $m(\zeta_{j},\vec{u})>1$. If the branch multiplicity at $\zeta_{j}$ is 2, then there are precisely two connected components of $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}\smallsetminus\operatorname{pr}^{-1}(\zeta_{j})$ containing the four roots of $f$, which violates the fact that the roots are equispaced. Thus $m(\zeta_{j},\vec{u})=4$. As $f$ has degree 4, this can only be true for one of the approximants to $V_{f,\infty}$, so we conclude that $\zeta_{j}=\zeta$ and $\vec{u}=\vec{w}$. Moreover, since the separable degree of the residue extension of $K[z]/(f)$ is 2, we conclude from (4.1) that $\deg_{\mathrm{s}}(\eta)=\deg_{\mathrm{s}}(\psi)=2$, where $\psi$ is the residual polynomial for $\phi_{j+1}$. It follows that $n=3$, that $j=2$, that $\phi_{2}$ is quadratic, and that $\phi_{j+1}=f$. In particular, $V_{2}/v$ either has ramification index 2 or purely inseparable residue extension of degree 2.
Now we know that $V_{3}=V_{f,\infty}$ and that the key polynomial $\phi_{2}$ for $V_{2}$ is quadratic. We take $\alpha$ to be a root of $\phi_{2}$. Let $r$ be the valuation of the difference between distinct roots of $f$. Using the notation of the previous paragraph, the fact that $\eta$ has separable degree 2 means that $\phi_{2}$ has roots in exactly two distinct maximal open subdisks of $D(\alpha,r)$. That is, the valuation of the difference between the roots of $\phi_{2}$ is $r$. To conclude, write $\phi_{2}(z)=z^{2}+bz+c$. The two roots of $\phi_{2}$ are $\alpha$ and $-\alpha-b$, so the valuation of the difference is
$r=v(2\alpha+b)\in v(K(\alpha)^{\times}),$
as desired. ∎
6 Elliptic Curves
We will give the proof of Theorem A in three parts:
  1. 1.
    Residue characteristic different from 2 or 3 in §6.1;
  2. 2.
    Residue characteristic different from 2 in §6.2; and
  3. 3.
    Residue characteristic 2 in §6.3.
The first part is completely elementary.
6.1 Residue characteristic different from 2 or 3
Theorem 6.1.
[…]
Proof.
Our hypothesis on the residue field ensures that $K$ itself has characteristic different from 2 or 3, and so we may assume that $E$ is given by a short Weierstrass equation of the form
$E\colon y^{2}=x^{3}+Ax+B,$
with $A,B\in K$. Define
$n=\min\left(3v(A),2v(B)\right).$
The discriminant $\Delta_{E}=-16(4A^{3}+27B^{2})$ satisfies $v(\Delta_{E})\geq n$. Define $L=K(\pi^{1/g})$, where $g=12/\gcd(n,12)$ and $\pi$ is a uniformizer for $K$. Note that $L/K$ is separable and totally ramified. Moreover, since $n$ either vanishes, or else is divisible by at least one of 2 or 3, we find that the degree $[L:K]$ is in the set $\{1,2,3,4,6\}$.
Set $u=\pi^{n/12}\in L$. Making the change of coordinates $x\mapsto u^{2}x$ and $y\mapsto u^{3}y$, we obtain a new curve
$E^{\prime}_{/L}\colon y^{2}=x^{3}+A^{\prime}x+B^{\prime},$
where
$v(A^{\prime})=v(A)-4v(u)=v(A)-\frac{n}{3}\geq 0\quad\text{and}\quad v(B^{\prime})=v(B)-6v(u)=v(B)-\frac{n}{2}\geq 0.$
Hence $E^{\prime}$ has coefficients in the valuation ring of $L$. The new discriminant satisfies
$v(\Delta_{E^{\prime}})=v(\Delta_{E})-12v(u)=v(\Delta_{E})-n\geq 0.$
If $v(\Delta_{E})=n$, then $E^{\prime}$ has good reduction. Otherwise, we find that $v(\Delta_{E})>n=3v(A)=2v(B)$. In particular, $v(A^{\prime})=0$, which implies that the reduction $\tilde{E}^{\prime}$ has a nodal singularity [20, III.1.4]. Since $n=3v(A)=2v(B)$, we find that $6\mid n$, and hence $g=[L:K]$ divides $2$. ∎
6.2 Residue characteristic different from 2
Throughout this entire section, we assume that $K$ has residue characteristic different from 2. Our goal is to prove the following:
Theorem 6.2.
Let $E_{/K}$ be an elliptic curve. There is a separable weakly totally ramified extension $L/K$ of degree dividing 12 such that $E$ admits a semistable model over $L$. If $E$ has potential multiplicative reduction, then we may take the degree of $L/K$ to divide 2.
The next lemma gives a criterion for determining reduction type of an elliptic curve in terms of the images of the 2-torsion in the residue field.
Lemma 6.3.
Let $E_{/K}$ be an elliptic curve given by a Weierstrass equation of the form
$E\colon y^{2}=(x-\alpha)(x-\beta)(x-\gamma),$
where $\alpha,\beta,\gamma\in K_{\mathrm{a}}$. Suppose that $\min\left\{v(\alpha),v(\beta),v(\gamma)\right\}\geq 0$.
  1. 1.
    If the reductions $\tilde{\alpha}$, $\tilde{\beta}$, and $\tilde{\gamma}$ are pairwise distinct, then $E$ has good reduction.
  2. 2.
    If exactly two of the reductions $\tilde{\alpha}$, $\tilde{\beta}$, and $\tilde{\gamma}$ coincide, then $E$ has multiplicative reduction.
  3. 3.
    If all three of the reductions $\tilde{\alpha}$, $\tilde{\beta}$, and $\tilde{\gamma}$ coincide, then $E$ has additive reduction.
In particular, this model of $E$ has semistable reduction if and only if $D(0,0)$ is the minimal disk containing $\alpha,\beta,\gamma$.
Remark 6.4.
The different reduction types correspond to different group structures on the special fiber of the Néron model $\mathcal{E}$. (In the setting of the lemma, it coincides with the locus of smooth points of $E$, viewed as a scheme over $\mathcal{O}_{K}$.) Assume the residue field of $K$ is algebraically closed for simplicity. If $E$ has good reduction, then the 2-torsion remains separate under reduction to yield the group
$\mathcal{E}_{s}[2]\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}.$
If $E$ has multiplicative reduction, then two of the 2-torsion points collapse to a singular point, and we have
$\mathcal{E}[2]\cong\mathbb{G}_{m}[2]\cong\mathbb{Z}/2\mathbb{Z}.$
If $E$ has additive reduction, then all of the nontrivial 2-torsion collapses to a singular point, and we have
$\mathcal{E}_{s}[2]\cong\mathbb{G}_{a}[2]=0.$
Proof of Lemma 6.3.
The discriminant of $E$ is given by
$\Delta_{E}=2^{4}(\alpha-\beta)^{2}(\alpha-\gamma)^{2}(\beta-\gamma)^{2}.$
If $\alpha$, $\beta$, $\gamma$ have distinct images in the residue field, then $\tilde{\Delta}_{E}=\Delta_{\tilde{E}}\neq 0$. Hence $E$ has good reduction.
If any of $\tilde{\alpha}$, $\tilde{\beta}$, and $\tilde{\gamma}$ coincide, then the formula in the last paragraph shows that $\tilde{\Delta}_{E}=\Delta_{\tilde{E}}=0$. So $E$ has bad reduction. Without loss of generality, we may assume that $\tilde{\alpha}=\tilde{\beta}$. If $\tilde{\gamma}=\tilde{\alpha}$, then the reduction of $E$ has the form $y^{2}=(x-\tilde{\alpha})^{3}$, which has a cuspidal singularity at $(\tilde{\alpha},0)$. That is, $E$ has additive reduction. If instead $\tilde{\gamma}\neq\tilde{\alpha}$, then $\tilde{\gamma}$ is a simple root of the reduction of the polynomial $(x-\alpha)(x-\beta)(x-\gamma)\in K[x]$. By Hensel’s Lemma, $\gamma\in K$. We make the change of coordinates $x\mapsto x+\gamma$ to get an equation of the form
$E^{\prime}\colon y^{2}=x(x-\delta)(x-\varepsilon)=x^{3}-(\delta+\varepsilon)x^{2}+\delta\varepsilon x,$
where $\tilde{\delta}=\tilde{\varepsilon}$. The reduction of this equation is
$y^{2}=(x-\tilde{\delta})^{3}+\tilde{\delta}(x-\tilde{\delta})^{2}.$
Since $\tilde{\delta}\neq 0$, the singularity at $(\tilde{\delta},0)$ is nodal, and $E$ has multiplicative reduction. ∎
Proposition 6.5.
Let $E_{/K}$ be an elliptic curve with potential multiplicative reduction. There is a $K$-rational $2$-torsion point $P$ such that in any model of $E$ over $\mathcal{O}_{K_{\mathrm{a}}}$ with multiplicative reduction, $P$ does not reduce to the singular point of the special fiber.
Proof.
Since $K$ does not have characteristic 2, we may begin by choosing a model for $E$ of the form $y^{2}=f(x)$ with $f\in K[x]$ a cubic polynomial. Let $\alpha,\beta,\gamma\in K_{\mathrm{a}}$ be the roots of $f$. We claim that there exists a relabeling of the roots of $f$ such that
$v(\beta-\gamma)>v(\alpha-\beta)=v(\alpha-\gamma).$
That is, $\beta$ and $\gamma$ are closer to each other than to $\alpha$. To see it, start with any labeling of the roots. Let $u\in K_{\mathrm{a}}$ be such that $v(u)=\min\Big{(}v(\alpha-\beta),v(\alpha-\gamma)\Big{)}$. Define
$g(x)=u^{-3}f(ux+\alpha)=x\left(x+\frac{\alpha-\beta}{u}\right)\left(x+\frac{\alpha-\gamma}{u}\right).$
The curve $E^{\prime}\colon y^{2}=g(x)$ is isomorphic to $E$ and satisfies the hypotheses of Lemma 6.3. As $E$ has potential multiplicative reduction and $\tilde{g}$ has at least two distinct roots, it must have exactly two distinct roots. If $\frac{\alpha-\beta}{u}$ and $\frac{\alpha-\gamma}{u}$ both have valuation zero, then we already have the correct labeling because $v(\alpha-\beta)=v(\alpha-\gamma)=v(u)$ and
$v(\beta-\gamma)=v\left(\frac{\alpha-\beta}{u}-\frac{\alpha-\gamma}{u}\right)+v(u)>v(u).$
Otherwise, exactly one of $v\Big{(}\frac{\alpha-\beta}{u}\Big{)}$ or $v\Big{(}\frac{\alpha-\gamma}{u}\Big{)}$ has positive valuation, say the former. So $v(\alpha-\gamma)=v(u)$, $v(\alpha-\beta)>v(u)$, and
$v(\beta-\gamma)=v\left(\frac{\alpha-\beta}{u}-\frac{\alpha-\gamma}{u}\right)+v(u)=v(\alpha-\gamma).$
Swapping $\alpha$ and $\gamma$ finishes the proof of the claim.
Next we claim that $f(x)$ has a $K$-rational root. Let $L$ be a splitting field for $f$. Every element of $G_{K}$ preserves the valuations $v(\alpha-\beta)$, $v(\alpha-\gamma)$, and $v(\beta-\gamma)$. But the third of these is strictly larger than the first two, so any element of $G_{K}$ fixes or swaps $\beta$ and $\gamma$. That is, every element of $G_{K}$ fixes $\alpha$. Since $\alpha$ is a root of the separable polynomial $f$, we conclude $\alpha$ is $K$-rational.
Let $P=(\alpha,0)\in E(K)$ be our distinguished 2-torsion point. On the model $E^{\prime}$ defined earlier in the proof, the point $P$ corresponds to $(0,0)$. By construction, $E^{\prime}$ has multiplicative reduction, and the double root of $\tilde{g}$ corresponds to the singular point of the reduced curve $\tilde{E}^{\prime}$. Consequently, $(0,0)$ does not reduce to the singular point. This property persists under any invertible change of coordinates over $\mathcal{O}_{K_{\mathrm{a}}}$. ∎
Proof of Theorem 6.2.
Choose a model for $E$ of the form $y^{2}=f(x)$ with $f\in K[x]$ a monic cubic polynomial. Write $\pi$ for a uniformizer of $K$. We consider four cases: the first three prove the general case of semistability, while the final case addresses the setting where $E$ has potential multiplicative reduction.
Case 1: $f$ admits a $K$-rational root. After an appropriate translation, we obtain a new model
$E^{\prime}_{/K}\colon y^{2}=xg(x),$
where $g$ is monic and quadratic. Let $m/n$ be the minimum valuation of a root of $g$, where $m\in\mathbb{Z}$ and $n\in\{1,2\}$. Set $L=K(\pi^{1/2n})$, and make the $L$-rational change of coordinates $(x,y)\mapsto(\pi^{m/n}x,\pi^{3m/2n}y)$ in order to obtain a third model
$E^{\prime\prime}_{/L}\colon y^{2}=xh(x).$
Now $h$ is monic with integral coefficients, and at least one of the roots of $h$ has valuation 0. Then $E^{\prime\prime}$ has semistable reduction (Lem. 6.3), and $L$ is totally ramified with degree dividing 4.
Case 2: $f$ is irreducible over $K$ and the extension $K[x]/(f)$ is weakly totally ramified. Let $\alpha$ be a root of $f$ in this extension. Replacing $K$ with the cubic extension $K(\alpha)$, we find ourselves in Case 1. That is, $E$ attains semistable reduction after a weakly totally ramified extension of degree dividing 12.
Case 3: $f$ is irreducible over $K$ and the extension $K[x]/(f)$ is unramified. Write $B$ for the minimal disk in $\mathbb{C}_{K}$ containing the roots of $f$; it contains a $K$-rational point (Lem. 2.6). Without loss of generality, we may translate by this point in order to assume that $B$ contains 0. Irreducibility of $f$ implies that all of its roots have the same valuation $r$, and the unramified hypothesis shows that $r\in\mathbb{Z}$. Set $L=K$ if $r$ is even and $L=K(\pi^{1/2})$ if $r$ is odd. Rescaling by the transformation $(x,y)\mapsto(\pi^{r}x,\pi^{3r/2}y)$ gives a new model
$E^{\prime}_{/L}\colon y^{2}=g(x),$
where $g(x)\in\mathcal{O}_{L}$ is monic, and all of its roots have valuation 0. The minimal disk about the roots of $g$ contains 0, and hence there exist roots $\alpha,\beta\in K_{{\mathrm{a}}}$ such that $v(\alpha-\beta)=0$. We conclude that $E^{\prime}$ has semistable reduction (Lem. 6.3), and that $L/K$ is totally ramified of degree dividing $2$.
Case 4: $E$ has potential multiplicative reduction. Let $P$ be a $K$-rational 2-torsion point as in Lemma 6.5. Without loss of generality, we may translate by the $x$-coordinate of this point in order to assume that our model has the form
$E\colon y^{2}=xg(x),$
where $g$ is a monic quadratic polynomial. Let $r$ be the minimum valuation of a root of $g$. We claim that $r\in\mathbb{Z}$. Assuming the claim for the moment, we define $L=K(\pi^{1/2})$ if $r$ is odd and $L=K$ if $r$ is even. Then $L/K$ is separable and weakly totally ramified of degree 1 or 2. Make the change of coordinates $(x,y)\mapsto(\pi^{r}x,\pi^{3r/2}y)$ to obtain an equation
$E^{\prime}_{/L}\colon y^{2}=xh(x),$
where $h$ is monic, and at least one of its roots has valuation 0. In fact, both roots have valuation 0, for otherwise $\tilde{E}^{\prime}$ has a singularity at $(0,0)$, which contradicts our choice of $P$. We conclude that $E^{\prime}$ has multiplicative reduction (Lem. 6.3).
Now we prove the claim that $r\in\mathbb{Z}$. Suppose not. Write $g(x)=x^{2}+bx+c$ with $b,c\in K$. The Newton polygon for $g$ must have a single segment with half-integral slope. It follows that $v(c)$ is odd, and $v(b)\geq\frac{1}{2}v(c)$. If $\gamma$ is a root of $g$, then $v(\gamma)=\frac{1}{2}v(c)$, and we can rescale the original model $E$ to get a new model of the form
$E^{\prime\prime}\colon y^{2}=xp(x),$
where $p(x)=\gamma^{-2}g(\gamma x)$. This model has semistable reduction since $xp(x)$ has at least two distinct roots over the residue field. But we know that $E^{\prime\prime}$ has multiplicative reduction, so we conclude that $p$ has a double root over the residue field. Hence, the valuation of its discriminant is positive:
$v\Big{(}(b^{2}-4c)\gamma^{-2}\Big{)}=v(b^{2}-4c)-2v(\gamma)=v(b^{2}-4c)-v(c)=v(b^{2}/c-4)>0.$
By the ultrametric inequality, we find that $v(b)=\frac{1}{2}v(c)$. But this is absurd since $v(c)$ is odd and $v(b)\in\mathbb{Z}$. The proof is complete. ∎
6.3 Residue characteristic 2
Throughout, assume that $K$ has residue characteristic 2. Our goal is to prove the following result:
Theorem 6.6.
Let $E_{/K}$ be an elliptic curve. There is a separable weakly totally ramified extension $L/K$ of degree dividing 24 such that $E$ has a model over $L$ with semistable reduction. If $E$ has potential multiplicative reduction, then we may take the degree of $L/K$ to divide 2.
An elliptic curve in this setting is given by the general Weierstrass model
$E_{/K}\colon y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6},$
where $a_{1},a_{2},a_{3},a_{4},a_{6}\in K$. We define the standard quantities
$b_{2}$ $=a_{1}^{2}+4a_{2}$
$b_{4}$ $=2a_{4}+a_{1}a_{3}$
$b_{6}$ $=a_{3}^{2}+4a_{6}$
$b_{8}$ $=a_{1}^{2}a_{6}+4a_{2}a_{6}-a_{1}a_{3}a_{4}+a_{2}a_{3}^{2}-a_{4}^{2}$
$\Delta$ $=-b_{2}^{2}b_{8}-8b_{4}^{3}-27b_{6}^{2}+9b_{2}b_{4}b_{6}$
$c_{4}$ $=b_{2}^{2}-24b_{4}.$
Recall that if all $a_{i}\in\mathcal{O}_{K}$, then $E$ has good reduction if $v(\Delta)=0$, and $E$ has multiplicative reduction if $v(\Delta)>0$ and $v(c_{4})=0$ [20, Prop. VII.5.1].
The 2-torsion of an elliptic curve over $K$ is ill-behaved under the reduction map, so we will want to focus our attention on 3-torsion points. The 3-division polynomial for $E$ is given by:
$\psi_{3}(x)=3x^{4}+b_{2}x^{3}+3b_{4}x^{2}+3b_{6}x+b_{8}.$ (6.1)
The roots of $\psi_{3}$ are precisely the $x$-coordinates of the nontrivial 3-torsion points for $E$. We can now state an analogue of Lemma 6.3 for 3-torsion.
Lemma 6.7.
Let $E_{/K}$ be an elliptic curve given by a Weierstrass model with coefficients in $\mathcal{O}_{K}$. Write $\psi_{3}\in\mathcal{O}_{K}[x]$ for the 3-division polynomial, and write $\tilde{\psi}_{3}$ for its image in $\tilde{K}[x]$.
  1. 1.
    If $\tilde{\psi}_{3}$ has four distinct roots in an algebraic closure, then $E$ has good reduction.
  2. 2.
    If $\tilde{\psi}_{3}$ has one simple root and one triple root in an algebraic closure, then $E$ has multiplicative reduction.
  3. 3.
    If $\tilde{\psi}_{3}$ has a quadruple root in an algebraic closure, then $E$ has additive reduction.
One of these three cases must occur. In particular, this model of $E$ has semistable reduction if and only if the minimal disk containing its roots is $D(0,0)$.
Proof.
The statement is geometric, so we may assume the residue field $\tilde{K}$ is algebraically closed. Let $\tilde{E}_{\mathrm{ns}}$ be the nonsingular locus of the reduction of $E$; it is a $\tilde{K}$-group scheme. Let us write $G=\tilde{E}_{\mathrm{ns}}[3]$ for the $3$-torsion subgroup. Observe that
$G\cong\begin{cases}\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}&\text{if$E$ has good reduction}\\ \mathbb{Z}/3\mathbb{Z}&\text{if $E$ has multiplicative reduction}\\ 0&\text{if $E$ has additive reduction}.\end{cases}$
If $\tilde{\psi}_{3}$ has four distinct roots, then at most one of them is the $x$-coordinate of a singular point. Each of the remaining roots gives rise to two nontrivial elements of $G$, so that $|G|\geq 6+1$. Hence $|G|=9$, and $E$ has good reduction.
Suppose now that $\tilde{\psi}_{3}$ has a multiple root, say $\tilde{x}_{0}$. We claim that $\tilde{x}_{0}$ is the $x$-coordinate of a singular point of $\tilde{E}$. One checks that $\frac{d\psi_{3}}{dx}=3\psi_{2}(x)$, where $\psi_{2}(x)=4x^{3}+b_{2}x^{2}+2b_{4}x+b_{6}$ is the $2$-division polynomial — the polynomial whose roots are the $x$-coordinates of 2-torsion points. It follows that $\tilde{x}_{0}$ satisfies $\psi_{2}$ as well. If $P=(\tilde{x}_{0},\tilde{y}_{0})$ were nonsingular on $\tilde{E}$, then $P$ would be both a 2-torsion point and a 3-torsion point, which is clearly impossible. In particular, since $\tilde{E}$ has at most one geometric singularity, we see that $\tilde{x}_{0}$ is the only multiple root of $\psi_{3}$.
If $\tilde{\psi}_{3}$ has a simple root and a triple root, then the triple root corresponds to a singularity on $\tilde{E}$, while the simple root gives rise to pair of nontrivial 3-torsion points in $G$. Thus $|G|=3$ and $E$ has multiplicative reduction.
If $\tilde{\psi}_{3}$ has two simple roots and a double root, then the two simple roots give rise to four nontrivial 3-torsion points in $G$. This implies $|G|\geq 5$, so that $|G|=9$ and $E$ has good reduction. But this is absurd since the double root of $\tilde{\psi}_{3}$ yields a singularity of the reduced curve $\tilde{E}$.
Finally, suppose that $\tilde{\psi}_{3}$ has a quadruple root. Then $\tilde{\psi}_{3}$ has no root corresponding to a nonsingular point of $\tilde{E}$ and $|G|=1$. Therefore, $E$ has additive reduction. ∎
Our next task is to determine how coordinate changes affect the 3-division polynomial. The most general coordinate change on $\mathbb{A}^{2}$ that yields a Weierstrass model for $E$ is
$(x,y)\mapsto(u^{2}x+r,u^{3}y+u^{2}sx+t),\qquad\text{where }r,s,t\in K,u\in K^{\times}.$
Note that this transformation can be written as a composition of four basic types of transformation:
  • homothety: $(x,y)\mapsto(u^{2}x,u^{3}y)$ for $u\in K^{\times}$;
  • $x$-translation: $(x,y)\mapsto(x+r,y)$ for $r\in K$;
  • shear: $(x,y)\mapsto(x,y+sx)$ for $s\in K$; and
  • $y$-translation: $(x,y)\mapsto(x,y+t)$ for $t\in K$.
The following lemma is proved by direct computation, or by staring at the formulas for how the $b_{i}$’s transform under a change of coordinates [20, p.45].
Lemma 6.8.
Let $E_{/K}$ be an elliptic curve in Weierstrass form:
$y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6}.$
The 3-division polynomial
$\psi_{3}(x)=3x^{4}+b_{2}x^{3}+3b_{4}x^{2}+3b_{6}x+b_{8}$
is unaffected by shears and $y$-translations. Homotheties have the effect $\psi_{3}(x)\mapsto u^{-8}\psi_{3}(u^{2}x)$, and $x$-translations have the effect $\psi_{3}(x)\mapsto\psi_{3}(x+r)$.
Proposition 6.9.
Let $E_{/K}$ be an elliptic curve given by a Weierstrass equation. Suppose that $E$ has potential multiplicative reduction. There is a $3$-torsion point $P=(\alpha,\beta)\in E(K_{\mathrm{a}})$ such that $\alpha\in K$, and in any model of $E_{K_{\mathrm{a}}}$ with multiplicative reduction, $P$ does not reduce to the singular point of the special fiber.
Proof.
Let $\psi_{3}$ be the $3$-division polynomial for $E$. Just as in the proof of Proposition 6.5, one shows that $\psi_{3}$ has a $K$-rational root $\alpha$ that is farther from the other three roots than they are from each other. ∎
Proof of Theorem 6.6.
Assume that $E$ is given by a general Weierstrass equation over $K$:
$E\colon y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6}.$
Let $\psi_{3}$ be the $3$-division polynomial for $E$, and let $\pi$ be a uniformizer for $K$.
Case 1: $E$ has a $K$-rational $3$-torsion point. We will show that $E$ has multiplicative reduction over $K$, or else it has a model with good reduction over $K(\pi^{1/3})$.
Writing $P=(x_{0},y_{0})$ for the $K$-rational 3-torsion point, we make the change of coordinates $(x,y)\mapsto(x+x_{0},y+y_{0})$ in order to assume that $P=(0,0)$, and consequently, $a_{6}=0$. Also, $a_{3}\neq 0$, for otherwise there would only be one $3$-torsion point with $x$-coordinate 00. We now make the change of coordinates $(x,y)\mapsto(x,x+a_{4}/a_{3}y)$ in order to adjust the tangent line at $(0,0)$ to be $y=0$. Since $(0,0)$ is a 3-torsion point, the line $y=0$ is a flex. Now we have the equation
$E^{\prime}_{/K}\colon y^{2}+axy+by=x^{3},$
with $a,b\in K$. The discriminant of the curve $E^{\prime}$ is
$\Delta^{\prime}=b^{3}(a^{3}-27b).$
If $v(a^{3})\leq v(b)$, we apply the rescaling $(x,y)\mapsto(a^{2}x,a^{3}y)$ to obtain the equation
$E^{\prime\prime}_{/K}:y^{2}+xy+(b/a^{3})y=x^{3}.$
This model has integral coefficients and $c_{4}$-invariant $1-24b/a^{3}$, which has valuation 0. That is, $E^{\prime\prime}$ has multiplicative reduction over $K$.
If $v(a^{3})>v(b)$, we write $m=v(b)$ and apply the rescaling $(x,y)\mapsto(\pi^{2m/3}x,\pi^{m}y)$ to obtain an integral equation with discriminant valuation 0. That is, we have a model with good reduction defined over the field $K(\pi^{1/3})$.
Case 2: $E$ has potential multiplicative reduction. Let $P=(x_{0},y_{0})$ be a 3-torsion point with $x_{0}\in K$ as in Proposition 6.9. Replacing $K$ with $K(y_{0})$, we are in Case 1 of the proof, and we conclude that $E$ has multiplicative reduction over $K(y_{0})$. If $K(y_{0})/K$ is weakly totally ramified, then set $L=K(y_{0})$. Otherwise, $K(y_{0})/K$ is unramified, and we set $L=K$. Since the reduction type does not change under unramified extension, it follows that $E$ has a model over $K$ with multiplicative reduction. In either case, the degree of $L/K$ divides 2.
Case 3: $\psi_{3}$ has a $K$-rational root. We claim that $E$ attains semistable reduction over a separable weakly totally ramified extension of degree dividing 6. Let $P=(x_{0},y_{0})$ be a $3$-torsion point on $E$ with $x_{0}\in K$. Set $K^{\prime}=K(y_{0})$. If $K^{\prime}/K$ is weakly totally ramified, then we find ourselves in Case 1 with $K$ replaced by $K^{\prime}$. If instead $K^{\prime}/K$ is unramified, then the proof of Case 1 shows that $E$ has a model over $K^{\prime}(\pi^{1/3})$ with semistable reduction. But the reduction type is unaffected by unramified extensions, so we also have semistable reduction over the field $K(\pi^{1/3})$.
Case 4: $\psi_{3}$ has a root defined over a weakly totally ramified extension of $K$. Let $K^{\prime}/K$ be a (separable) weakly totally ramified extension that contains a root of $\psi_{3}$. We may assume that the degree $[K^{\prime}:K]$ divides 4. Replacing $K$ with $K^{\prime}$, we find ourselves in Case 3, and hence $E$ attains semistable reduction over a separable weakly totally ramified extension of degree dividing 24.
Case 5: $\psi_{3}$ has no root defined over a weakly totally ramified extension of $K$. In particular, $E$ has potential good reduction (Prop. 6.9), so the roots of $\psi_{3}$ are equispaced (Prop. 6.7). Let $B$ be the minimal disk about the roots of $\psi_{3}$. There is $r\in B$ such that $K(r)/K$ is separable and weakly totally ramified of degree dividing 2, and such that $v(\alpha-\beta)\in v(K(r)^{\times})$ for every pair of distinct roots $\alpha,\beta$ of $\psi_{3}$ (Lem. 5.5). Let $m$ be this common value, so that $B=D(\alpha,m)$. Let $u\in K_{\mathrm{a}}$ be such that $u^{2}\in K(r)$ and $v(u^{2})=m$. Making the change of coordinates $(x,y)\mapsto(u^{2}x+r,u^{3}y)$, we obtain a new model
$E^{\prime}_{/K^{\prime}}\colon y^{2}+a_{1}^{\prime}xy+a_{3}^{\prime}y=x^{3}+a_{2}^{\prime}x^{2}+a_{4}^{\prime}x+a_{6}^{\prime},$
where $K^{\prime}=K(u,r)$. Note $[K^{\prime}:K]$ divides 4. By construction, the minimal disk about the roots of the 3-division polynomial $\psi_{3,E^{\prime}}$ is $D(0,0)$. We now have two subcases; the argument in each case is essentially the same, though the details are quite different.
Case 5a: $K$ has characteristic 0. Without loss of generality, we may assume that $a_{1}^{\prime}=a_{3}^{\prime}=0$ after making the change of coordinates $(x,y)\mapsto(x,y-a_{1}^{\prime}/2x-a_{3}^{\prime}/2)$. Since this is a composition of a shear and a $y$-translation, it does not affect the 3-division polynomial (Lem. 6.8). We have
$\psi_{3,E^{\prime}}(x)=3x^{4}+4a_{2}^{\prime}x^{3}+6a_{4}^{\prime}x^{2}+12a_{6}^{\prime}x+\left(4a_{2}^{\prime}a_{6}^{\prime}-(a_{4}^{\prime})^{2}\right).$
Define
$A=4a_{2}^{\prime},\quad B=2a_{4}^{\prime},\quad C=4a_{6}^{\prime}.$
Then $A,B,C\in\mathcal{O}_{K^{\prime}}$ since we have arranged for $\psi_{3,E^{\prime}}$ to have integral coefficients. Also, $\psi_{3,E^{\prime}}$ has four distinct roots over $k_{{\mathrm{a}}}$, so at least one of $A,C$ has valuation 0. Moreover, the constant coefficient of $\psi_{3,E^{\prime}}$ must have valuation 0. For if not, $\tilde{\psi}_{3,E^{\prime}}$ would have 00 as a simple root, which would lift to a root of $\psi_{3,E^{\prime}}$ in $K^{\prime}$, contradicting the fact that $\psi_{3,E}$ does not have a root defined over a weakly totally ramified extension of $K$. It follows that
$v(AC-B^{2})=v\left(16a_{2}^{\prime}a_{6}^{\prime}-4(a_{4}^{\prime})^{2}\right)=2v(2).$
Suppose first that $A$ has valuation 0. Since $v(a_{2}^{\prime})=-2v(2)$, the minimal disk about the roots of $z^{2}-a_{2}^{\prime}$ is $D:=D(\sqrt{a_{2}^{\prime}},0)$. We use Ramified Approximation to choose $s\in D$ such that $K^{\prime}(s)/K^{\prime}$ is weakly totally ramified and at worst quadratic. Next, define $t=a_{4}^{\prime}/(2s)$. Note that $v(s)=-v(2)$ and $v(t)=v(a_{4}^{\prime})\geq-v(2)$. Set $L=K(s)$. Making the change of coordinates $(x,y)\mapsto(x,y+sx+t)$, we obtain the new model
$E^{\prime\prime}_{/L}\colon y^{2}+a_{1}^{\prime\prime}xy+a_{3}^{\prime\prime}=x^{3}+a_{2}^{\prime\prime}x^{2}+a_{6}^{\prime\prime},$
where
$a_{1}^{\prime\prime}=2s,\qquad a_{2}^{\prime\prime}=a_{2}^{\prime}-s^{2},\qquad a_{3}^{\prime\prime}=2t,\qquad a_{6}^{\prime\prime}=a_{6}^{\prime}-t^{2}.$
Evidently, $v(a_{1}^{\prime\prime})=0$ and $v(a_{3}^{\prime\prime})\geq 0$. If we write $s=\sqrt{a_{2}^{\prime}}+u$ with $v(u)\geq 0$, then $s^{2}=a_{2}^{\prime}+u\sqrt{A}+u^{2}$. In particular, we find $v(a_{2}^{\prime\prime})\geq 0$ and
$v(a_{6}^{\prime\prime})=v\left(\frac{AC-B^{2}}{16s^{2}}+\frac{C(u\sqrt{A}+u^{2})}{4s^{2}}\right).$
Both displayed terms have nonnegative valuation, so $v(a_{6}^{\prime\prime})\geq 0$. Thus, $E^{\prime\prime}$ is an integral model, and it has good reduction since we haven’t changed the $3$-division polynomial (Lem. 6.7). The extension $L/K$ is weakly totally ramified of degree dividing 8.
If instead $C$ has valuation 0, then the minimal disk about the roots of $z^{2}-a_{6}^{\prime}$ is $D(\sqrt{a_{6}^{\prime}},0)$. We choose $t$ inside this disk with the Approximation Theorem, and we set $s=a_{4}^{\prime}/(2t)$. After the change of coordinates $(x,y)\mapsto(x,y+sx+t)$, the argument is similar to the previous paragraph.
Case 5b: $K$ has characteristic 2. The 3-division polynomial is given by
$\psi_{3,E^{\prime}}(x)=x^{4}+(a_{1}^{\prime})^{2}x^{3}+a_{1}^{\prime}a_{3}^{\prime}x^{2}+(a_{3}^{\prime})^{2}x+b_{8}^{\prime}.$
The reduction $\tilde{\psi}_{3,E^{\prime}}$ has four distinct roots, so at least one of $a_{1}^{\prime},a_{3}^{\prime}$ has valuation 0. Also, $v(b_{8}^{\prime})=0$ since otherwise $\psi_{3,E}$ has a $K^{\prime}$-rational root (cf. Case 5a).
Suppose first that $v(a_{1}^{\prime})=0$. The roots of $z^{2}+a_{1}^{\prime}z+a_{2}^{\prime}$ are $\alpha$ and $\alpha+a_{1}^{\prime}$, so that the valuation of their difference is 0. That is, the minimal disk about the roots of this polynomial is $D:=D(\alpha,0)$. Apply Ramified Approximation to obtain $s\in D$ with $K^{\prime}(s)/K^{\prime}$ weakly totally ramified and at worst quadratic. Set $t=(a_{4}^{\prime}+sa_{3}^{\prime})/a_{1}^{\prime}$. Set $L=K^{\prime}(s)$. Making the change of coordinates $(x,y)\mapsto(x,y+sx+t)$ gives the new model
$E^{\prime\prime}_{/L}\colon y^{2}+a_{1}^{\prime}xy+a_{3}^{\prime}y=x^{3}+a_{2}^{\prime\prime}x^{2}+a_{6}^{\prime\prime},$
where
$a_{2}^{\prime\prime}=s^{2}+a_{1}^{\prime}s+a_{2}^{\prime},\qquad a_{6}^{\prime\prime}=a_{6}^{\prime}+ta_{3}^{\prime}+t^{2}.$
Writing $s=\alpha+u$ for some $u$ with $v(u)\geq 0$, we see that $v(a_{2}^{\prime\prime})=u(u+a_{1}^{\prime})$, so that $v(a_{2}^{\prime\prime})\geq 0$. Since this change of coordinates is a composition of a shear and a $y$-translation, we find $b_{8}^{\prime}=b_{8}^{\prime\prime}$ (Lem. 6.8). Hence,
$0=v(b_{8}^{\prime\prime})=v\left((a_{1}^{\prime})^{2}a_{6}^{\prime\prime}+a_{2}^{\prime\prime}(a_{3}^{\prime})^{2}\right).$
Since $a_{2}^{\prime\prime}(a_{3}^{\prime})^{2}$ is integral and $a_{1}^{\prime}$ has valuation 0, we conclude that $a_{6}^{\prime\prime}$ is integral. That is, the model $E^{\prime\prime}$ is integral, and it has good reduction since we haven’t changed the $3$-division polynomial (Lem. 6.7). We also see that $[L:K]$ divides 8, and $L/K$ is weakly totally ramified.
If instead $v(a_{3}^{\prime})=0$, then the minimal disk about the roots of $z^{2}+a_{3}^{\prime}z+a_{6}^{\prime}$ is $D(\beta,0)$ for $\beta$ a root. We choose $t$ inside this disk with Ramified Approximation, and we set $s=(a_{4}^{\prime}+ta_{1}^{\prime})/a_{3}^{\prime}$. After the change of coordinates $(x,y)\mapsto(x,y+sx+t)$, the argument mimics the previous paragraph. ∎
7 Dynamical systems on $\mathbb{P}^{1}$
Let $f\in K(z)$ be a rational function of degree $d\geq 2$. Write $f=f_{0}/f_{1}$, where $f_{0},f_{1}$ are coprime polynomials with integral coefficients. After rescaling by a common element of $K^{\times}$, we may assume that some coefficient of $f_{0}$ or $f_{1}$ lies in $\mathcal{O}_{K}^{\times}$. Let $F_{0},F_{1}\in\mathcal{O}_{K}[X,Y]$ be homogeneous polynomials of degree $d$ such that $f_{i}=F_{i}(z,1)$. Write $\mathrm{Res}(F_{0},F_{1})$ for the resultant of $F_{0},F_{1}$; it is a homogeneous polynomial of degree $22d+$ 2 in the coefficients of $F_{0},F_{1}$ [19, §2.4]. Following Rumely, we write $\mathrm{ordRes}(f)$ for $v(\mathrm{Res}(F_{0},F_{1}))$. This quantity is independent of the choice of $F_{0},F_{1}$, and necessarily $\mathrm{ordRes}(f)\geq 0$. We say $f$ has semistable reduction if
$\mathrm{ordRes}(f)=\min_{\sigma\in\operatorname{PGL}_{2}(\mathbb{C}_{K})}\mathrm{ordRes}\left(\sigma^{-1}\circ f\circ\sigma\right).$
We say $f$ has good reduction if $\mathrm{ordRes}(f)=0$.
The Gauss point of $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ is the type II point $\zeta_{0,0}$. Given any other type II point $\zeta\in\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$, there is an element $\sigma\in\operatorname{PGL}_{2}(\mathbb{C}_{K})$ such that $\sigma(\zeta_{0,0})=\zeta$, and this $\sigma$ is unique up to left-multiplication by elements of $\operatorname{PGL}_{2}(\mathcal{O}_{\mathbb{C}_{K}})$. Define $f^{\sigma}=\sigma^{-1}\circ f\circ\sigma$. Then we obtain a function $\zeta\mapsto\mathrm{ordRes}(f^{\sigma})$ that does not depend on the choice of $\sigma$, and we can extend it to a continuous piecewise affine map $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}\to\mathbb{R}\cup\{\infty\}$. Rumely showed that this map is convex up on paths in $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$, and he defined the minimal resultant locus — denoted $\mathrm{MinResLoc}(f)$ — to be the subset of $\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}$ on which the function $\zeta\mapsto\mathrm{ordRes}(f^{\sigma})$ attains its minimum value [15, Thm. 1.1]. The type II points of the minimal resultant locus correspond to coordinate changes $f^{\sigma}$ that are semistable [16, Thm. C].
Proof of Theorem B.
Suppose first that $f$ has potential good reduction. In [3, §5], Benedetto explains how to achieve the desired result: replace Ax’s Lemma with a refinement that controls ramification. Corollary 1.2 is that refinement, and the result follows from Benedetto’s discussion.
Now suppose that $f$ is a general rational function of degree $d\geq 2$ on $\mathbb{P}^{1}_{K}$. We follow Rumely’s argument in [15, Thm. 3.2], making the necessary modifications. Define $a=f(\infty)$, and let $\mathrm{Fix}(f)$ be the set of type I fixed points of $f$ in $\mathbf{P}^{1,\mathrm{an}}_{\mathbb{C}_{K}}=\mathbf{A}^{1,\mathrm{an}}_{\mathbb{C}_{K}}\cup\{\infty\}$. The minimal resultant locus is contained in the convex hull of $\mathrm{Fix}(f)\cup f^{-1}(a)$, and it is either a single type II point or a segment with type II endpoints [15, Thm. 1.1]. Since $f$ is defined over $K$, the set $\mathrm{MinResLoc}(f)$ is stable under the action of $G_{K}$, though not necessarily pointwise fixed.
Let $Q\in\mathrm{MinResLoc}(f)$ be a point that is fixed by $G_{K}$ [15, Thm. 3.4]. Then $Q$ corresponds to a $G_{K}$-invariant disk $D$. Since $\mathrm{MinResLoc}(f)$ is contained in the convex hull of $\mathrm{Fix}(f)\cup f^{-1}(a)$, there exists $P\in\mathrm{Fix}(f)\cup(f^{-1}(a)\smallsetminus\{\infty\})$ such that $Q$ lies on the segment $[P,\infty]$. That is, $D$ contains $P$. If $P\in\mathrm{Fix}(f)$, then $P$ has degree at most $d+1$, while if $P\in f^{-1}(a)\smallsetminus\{\infty\}$, then $P$ has degree at most $d-1$. Since $D$ is $G_{K}$-invariant, every conjugate of $P$ lies in $D$. By Corollary 1.2, there is an element $\alpha\in K_{\mathrm{a}}\cap D$ such that $K(\alpha)/K$ is separable and weakly totally ramified, and $[K(\alpha):K]\leq\max\{q(d+1),q(d-1)\}$.
Now $Q$ lies on the segment $[\alpha,\infty]$. Let $Q_{0}\in\mathrm{MinResLoc}(f)$ be the closest point to $\alpha$; i.e., $[\alpha,Q_{0}]\cap\mathrm{MinResLoc}(f)=\{Q_{0}\}$. The function $\mathrm{ordRes}(\cdot)$, when restricted to $[\alpha,\infty]$, has a break in linearity at $Q_{0}$. The discussion after (2.9) in [15] shows there are integers $m,n$ with $11\leq n\leq d+$ 1 such that $Q_{0}=\zeta_{\alpha,m/n}$. (N.B. — We are working with valuations, whereas Rumely works with absolute values.) Taking a uniformizer $\pi$ for $K(\alpha)$ and an element $c\in K$ with $v(c)$ sufficiently large, we find that $z^{n}+cz+\pi^{m}$ is separable and has a root $\beta$ with valuation $m/n$. It follows that $Q_{0}$ is defined over $K(\alpha,\beta)$, which is separable and weakly totally ramified of degree at most $(d+1)\ \max\{q(d+1),q(d-1)\}$. Now conjugate $f$ by $z\mapsto\beta z+\alpha$ in order to obtain a model with minimal resultant, and hence semistable reduction. ∎